### v1

# When Hyperspectral Image Classification Meets Diffusion Models: An Unsupervised Feature Learning Framework

Abstract
Learning effective spectral-spatial features is important for the hyperspectral image (HSI) classification task, but the majority of existing HSI classification methods still suffer from modeling complex spectral-spatial relations and characterizing low-level details and high-level semantics comprehensively. As a new class of record-breaking generative models, diffusion models are capable of modeling complex relations for understanding inputs well as learning both high-level and low-level visual features. Meanwhile, diffusion models can capture more abundant features by taking advantage of the extra and unique dimension of timestep $t$. In view of these, we propose an unsupervised spectral-spatial feature learning framework based on the diffusion model for HSI classification for the first time, named Diff-HSI. Specifically, we first pretrain the diffusion model with unlabeled HSI patches for unsupervised feature learning, and then exploit intermediate hierarchical features from different timesteps for classification. For better using the abundant timestep-wise features, we design a timestep-wise feature bank and a dynamic feature fusion module to construct timestep-wise features, adaptively learning informative multi-timestep representations. Finally, an ensemble of linear classifiers is applied to perform HSI classification. Extensive experiments are conducted on three public HSI datasets, and our results demonstrate that Diff-HSI outperforms state-of-the-art supervised and unsupervised methods for HSI classification.
Index Terms:
Hyperspectral image classification, denoising diffusion probabilistic model, unsupervised feature learning.
I Introduction
HYPERSPECTRAL image (HSI) classification plays a crucial role in remote sensing [1, 2, 3, 4], as it aims to distinguish each pixel’s category in hyperspectral data by using dense and detailed electromagnetic spectral information [5]. The main objective is to effectively discriminate objects of interest by capturing subtle variations in the spectral signatures associated with their pixels [6]. To this end, a large number of methods are developed, relying on hand-crafted features [7, 8, 9, 10, 11]. However, these works are designed to extract features manually, leading to limited discriminative information extraction from the high spectral variability [12].
The rapid advancement of deep learning has led to the emergence of supervised approaches, a typical category of HSI classification methods, specifically utilizing deep convolutional neural networks (CNNs) [13, 14, 15] and transformers [16, 17, 18, 19]. Although these previous supervised methods have already achieved promising results in HSI classification, they still have limitations in spectral-spatial feature extraction. Firstly, previous supervised methods train the CNN or transformer network to learn explicit mappings which leads to poor performance to model complex spectral-spatial relations. Secondly, supervised methods tend to prioritize features from labeled areas while HSIs have large unlabeled regions that contain label-agnostic information. They result in a preference for learning labeling priors and limited general representation generation that characterize spectral-spatial relations. Additionally, the limited availability of labeled samples in HSI datasets poses constraints on previous supervised methods.
((a))
((b))
Figure 1: Different unsupervised feature learning frameworks including (a) existing unsupervised feature learning and (b) ours proposed Diff-HSI that extract timestep-wise features and construct multi-timestep representations by dynamic fusion for the first time.
As another category of HSI classification methods, [20, 21, 22, 23] is designed to learn spectral-spatial features from HSIs in an unsupervised manner. Typically, these methods leverage an encoder-decoder network trained in an unsupervised manner for HSI reconstruction and extract the spectral-spatial feature from the intermediate feature maps, depicted in Fig. 1(a). Although these methods mine implicit spectral attributes and spatial relations through unsupervised feature learning, the reconstruction networks they designed mainly capture low-level reconstruction features such as texture and edge, lacking high-level features which are necessary for discriminative representations to model the spectral-spatial relations comprehensively [24].
Learning rich representations and modeling the complex spectral-spatial relations from high dimensions are significant and should be considered for HSI classification. Most recently, diffusion models [25, 26, 27] have emerged as new generative models with superior performance in generation and reconstruction tasks, and have started to be explored in other computer vision tasks like semantic segmentation [28, 29, 30]. Diffusion models have the ability in the reconstruction task, so they are well-suited as a framework for unsupervised feature learning. Furthermore, there exist two appealing properties of using diffusion models [31, 32, 33] for unsupervised feature learning of hyperspectral data: 1) diffusion models learn implicit mappings, which have superior representation ability on complex problems via simple expressions and extract both high-level and low-level features from images; 2) the stepwise reverse denoising process of diffusion models are modeled as an iterative optimization process optimized by Langevin dynamics. This process brings a large number of degrees of freedom and tries to predict further information conditioned on the given noise-destructed data at every timestep, leading to better generalization to complex problems. When applying diffusion models to unsupervised feature learning of HSI data, it not only learns both high-level and low-level features but also extracts spectral-spatial features in the dimension of timestep $t$, as shown in Fig. 1(b).
Naturally, it is crucial to explore how to exploit abundant timestep-wise features more effectively and efficiently. On the one hand, the timestep-wise features are diverse and informative and may focus on different information. The shallow timestep-wise features are more informative for fine details, while deeper ones are more concerned about high-level semantics and global information [25, 34]. On the other hand, HSIs from various datasets acquired by different sensors exhibit distinct spectral characteristics, leading to variations in the spectral representation patterns of timestep-wise features extracted from different HSI datasets. Meanwhile, the abundant timestep-wise features pose a challenge in terms of computation overheads. In addition, manually selecting timestep-wise features to characterize spectral-spatial features effectively is struggle. Therefore, a dynamic feature fusion module that adaptively fuses features of different timesteps for different HSI datasets is desired.
In view of these, we propose the unsupervised spectral-spatial feature learning framework based on the diffusion model for HSI classification, named Diff-HSI. More specifically, the proposed framework first pretrain a diffusion model with unlabeled HSI patches for unsupervised feature learning, which is able to automatically learn label-agnostic knowledge and mine the connotation of unlabeled data that reveals the complex spectral-spatial dependencies compared to conventional supervised methods. Then, we extract features from different hierarchies from the pretrained denoising U-Net decoder to construct the timestep-wise feature bank, which contains various informative features from different timestep $t$. To effectively harness the features from the timestep-wise feature bank and softly learn the proper timestep-wise feature combination for different HSI datasets, we propose a dynamic feature fusion module. This module is designed to adaptively fuse hierarchical features from different timesteps, generating multi-timestep representations enriched with comprehensive spectral-spatial information. Ultimately, an ensemble of linear classifiers is employed to leverage these dynamic representations for accurate HSI classification.
To summarize, our contributions are listed as follows.
  • 1)
    As a pioneer work, we introduce diffusion models to HSI classification and propose a diffusion-based unsupervised feature learning framework called Diff-HSI, which can implicitly learn both high-level and low-level features to model complex spectral-spatial relations.
  • 2)
    To leverage abundant timestep-wise features effectively and efficiently, we construct a timestep-wise feature bank with hierarchical features from different timesteps, and design a dynamic feature fusion module to fuse these timestep-wise features adaptively. The above process learns the soft timestep-wise feature combination for each labeled patch of different HSI datasets and constructs comprehensive multi-timestep representations.
  • 3)
    Compared with several state-of-the-art supervised and unsupervised methods, experimental results demonstrate that our proposed method achieves superior classification accuracy on three public HSI datasets.
The remainder of this paper is organized as follows. Section II describes related work. In Section III, our proposed Diff-HSI is introduced in detail. Section IV conducts extensive experiments on three HSI datasets to demonstrate the effectiveness of the proposed method. Finally, some conclusions are drawn in Section V.
II Related Work
II-A Supervised-based HSI Classification Methods
Due to the remarkable breakthroughs achieved by deep learning in various computer vision tasks, many progressive deep learning-based networks have been widely utilized for supervised HSI classification methods, e.g., autoencoders (AEs) [21], deep belief networks (DBNs) [35] and recurrent neural networks (RNNs) [36]. Among these, CNNs draw significant attention with their excellent feature extraction capability and become mainstream in HSI classification [13, 14, 15]. CNNs can effectively extract spectral-spatial features through their ability to extract spatially structural information and locally contextual information. As the vision transformer rises in the field of computer vision, researchers begin to explore the value of the transformer in HSI classification and find it outperforms CNN-based methods in solving the problem of long-term spectral information dependencies [16, 17, 18, 19]. Although these supervised methods have achieved promising performance, they still require large amounts of labeled HSI samples, and are limited to modeling complex relations and capturing label-agnostic information.
II-B Unsupervised-based HSI Classification Methods
Unsupervised feature learning is a feature extraction paradigm to automatically learn feature representations from the input data without any annotated information. Leveraging unsupervised feature learning on unlabeled HSI patches is a solution to the limited labeled samples of HSI datasets. Typically, the commonly used unsupervised feature learning methods in HSI classification are based on the encoder–decoder paradigm, where an autoencoder-like network encodes the input HSI patches into a refined feature and then reconstructs the feature to initial HSI data by a decoder. Mou et al. [20] first design a fully 2D Conv–Deconv network in an end-to-end manner for unsupervised feature learning of HSI classification. Similarly, Mei et al. [21] design a 3D convolutional autoencoder (3D-CAE) for unsupervised feature learning of HSI classification. To alleviate the insufficiency of geometric representation and exploit the multi-scale features, Zhang et al. [22] design a multi-scale CNN-based unsupervised feature learning framework, with two branches of decoder and clustering optimized by the error feedback of image reconstruction and pseudo-label classification. In the aforementioned approaches, the feature representations learned from the unlabeled HSI patches contain more spectral-spatial information than supervised methods; however, they mainly focus on low-level visual concepts and ignore high-level feature extraction for HSI classification, which is solved in our methods.
II-C Diffusion Models
Diffusion models are a class of probabilistic generative models that progressively inject a standard Gaussian noise, then learn a model to reverse this process for sample generation [25, 26, 27]. Current research on diffusion models is mostly based on three formulations: denoising diffusion probabilistic models (DDPMs) [25], score-based generative models [33], and stochastic differential equations [37]. DDPMs are the mainstream diffusion models, and a large number of recent works based on DDPMs have made DDPMs increasingly powerful in terms of generative quality and diversity over other generative models [38]. Meanwhile, DDPMs have been widely used in several applications, including super-resolution [39], inpainting [40], and point cloud generation [41]. However, to the best of our knowledge, few works successfully adopt them to HSI classification. When applying diffusion models to HSI classification, it can not only implicitly learn both high-level and low-level features but also extract spectral-spatial features in the dimension of timestep $t$, leading to improvement in modeling complex spectral-spatial relations.
Figure 2: Overview of our proposed Diff-HSI. The method consists of two steps. Step 1: We pretrain the DDPM with HSI patches in an unsupervised manner and then freeze the parameters. Step 2: We construct a timestep-wise feature bank by extracting hierarchical features from the fixed pretrained U-Net decoder in terms of different timestep $t$. In order to softly fuse the features, a dynamic feature fusion module is designed to obtain effective multi-timestep representations adaptively. Finally, an ensemble of lightweight classifiers is trained by leveraging the multi-timestep representations of label-limited HSI patches to perform HSI classification.
III Method
Our proposed Diff-HSI aims to learn effective spectral-spatial representations in an unsupervised manner via denoising diffusion probabilistic model (DDPM) for HSI classification. The framework is shown in Fig. 2. Our method consists of two key steps. The first step is to pretrain the DDPM with unlabeled patches to learn complex spectral-spatial distributions and dependencies of HSI data. After pretraining, we freeze the parameters of the pretrained DDPM. The second step of our Diff-HSI is a supervised learning process aiming to fully extract and dynamically fuse the timestep-wise features from the pretrained denoising U-Net for classification. In order to capture the characteristics of different objects in HSI data, we first extract abundant spectral-spatial features from different timestep $t$ to construct the timestep-wise feature bank. Then, we design a dynamic feature fusion module that fuses the features in the feature bank dynamically and generates multi-timestep representations with discriminative information for classification. Finally, the representations are fed into an ensemble of lightweight classifiers to predict the label of the center pixel. In this section, we introduce our proposed Diff-HSI in detail, including the unsupervised diffusion-based pretraining step as well as the supervised finetuning step with multi-timestep representation generation.
III-A Diffusion-based Unsupervised Spectral-Spatial Feature Learning
In order to learn complex spectral-spatial relations and label-agnostic information of HSI data, we pretrain a diffusion model in an unsupervised manner in the first step of our Diff-HSI. We first give a brief review of DDPM and introduce the basic principles to make our method easier to understand. Then, we explain our unsupervised feature learning procedure in detail through diffusion-based pretraining with HSI data.
III-A1 A Brief Review of DDPM
Denoising diffusion probabilistic models are a class of likelihood-based models that reconstruct the distribution of training data via an encoder-decoder denoising model. The denoising model is trained to remove noise from the training data destructed by Gaussian noises step-by-step. These models consist of a forward noising process and a reverse denoising process. In the forward process, Gaussian noise is added to the original training data $x_{0}\sim q\left(x_{0}\right)$ step by step over $T$ time steps, which follows the Markovian process:
$q(x_{t}|x_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (1)
where $\mathcal{N}(.)$ is a Gaussian distribution, and the Gaussian variances $\{\beta_{t}\}_{t=0}^{T}$ that determines the noise schedule are either be learned or scheduled. The above formulation leads that an arbitrary noisy sample $x_{t}$ for each timestep $t$ is obtained directly from $x_{0}$:
$x_{t}=\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{(1-\overline{\alpha}_{t})}\epsilon,\epsilon\sim\mathcal{N}(0,I)$ (2)
where $\alpha_{t}=1-\beta_{t}$, and $\overline{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. Then in the reverse process, DDPM also follows a Markovian process to denoise the noisy sample $x_{T}$ to $x_{0}$ step by step. Under large $T$ and small $\beta_{t}$, the reverse transitions probability is approximated as a Gaussian distribution and is predicted by a learned neural network as follows:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}(x_{t},t))$ (3)
where the reverse process is re-parameterized by estimating $\mu_{\theta}(x_{t},t)$ and $\sigma_{\theta}(x_{t},t)$. $\sigma_{\theta}(x_{t},t)$ is set to $\sigma_{t}^{2}\mathbf{I}$, where $\sigma_{t}^{2}$ is not learned. In practice, rather than predicting $\mu_{\theta}(x_{t},t)$ directly, predicting the noise $\epsilon$ in Eq. 2 via a U-Net works best, and the parameterization of $\mu_{\theta}(x_{t},t)$ is derived as follows:
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$ (4)
The U-Net denoising model $\epsilon_{\theta}(x_{t},t)$ is optimized by minimizing the following loss function:
$\mathcal{L}(\theta)=E_{t,x_{0},\epsilon}[(\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0})+\sqrt{1-\overline{\alpha}_{t}\epsilon},t)^{2}]$ (5)
In our work, improved DDPM [27] is adopted and has been proven to bring some improvements to the above DDPM. In detail, learned variances $\sigma_{\theta}(x_{t},t)$ and an improved cosine noise schedule proposed in [27] lead to enhanced distribution learning ability.
Figure 3: 2D t-SNE visualization of features from different timesteps on the Indian Pines dataset. (a) original spectral feature. (b) feature at $t=100$. (c) feature at $t=300$. (d) feature at $t=500$. (e) feature at $t=700$. (f) feature at $t=900$.
III-A2 Unsupervised Hyperspectral Diffusion Pretraining
Using the training skills and optimization objectives mentioned above, our DDPM is trained by unlabeled hyperspectral data. Suppose that $S$ is the set of all samples in a dataset, $D$ is a denoising diffusion probabilistic model, and $E$ is an ensemble of lightweight classifiers. The target of the unsupervised pretraining step is to learn label-agnostic spectral-spatial features on unlabelled data sampled from $S$ so that the features can be leveraged in the finetuning step to train ensemble lightweight classifiers using a small amount of labeled data sampled from $S_{labeled}\in S$ for classification.
In order to learn effective spectral-spatial features in an unsupervised manner, we first pretrain a denoising diffusion probabilistic model with unlabeled patches randomly cropped from the HSI dataset. Before training, the data is pre-processed by principal components analysis (PCA) and patch cropping operation. Then, patches are randomly sampled from compressed HSI for DDPM pretraining. In detail, given an unlabeled patch $x_{0}\in\mathcal{R}^{H\times H}$, we gradually add Gaussian noise to the HSI patch according to the variance schedule $\{\beta_{t}\}_{t=0}^{T}$ in the diffusion process where $T$ is the total number of the timestep. Then, in the reverse process, a denoising U-Net is trained to predict the noise added on $x_{t-1}$ taking noisy patch $x_{t}$ and timestep $t$ as inputs. And $x_{t-1}$ is calculated by subtracting the predicted noise from $x_{t}$. In other words, the model tries to capture some useful information of $x_{t}$ to reconstruct $x_{t-1}$ that is closer to $x_{0}$. In each step of the training process, the timestep $t$ is randomly sampled from 0 to $T$.
In the pretraining procedure with unlabeled hyperspectral data, the DDPM is able to automatically learn the label-agnostic prior of HSI data that reveals the complex dependencies and relations in the spectral and spatial dimensions. At different timesteps, different input noisy patches and their optimization goals allow the DDPM to attend to varied information of the data, which is demonstrated to be critical to HSI classification.
III-B Multi-Timestep Representation for Supervised Finetuning
After unsupervised pretraining, in order to explore the rich information learned by the pretrained DDPM, the second step of our method is to extract effective features from the pretrained DDPM and construct multi-timestep representations for HSI classification. We first extract abundant features and construct a timestep-wise feature bank. Then, a dynamic feature fusion module is designed to automatically fuse the features into multi-timestep representations. Finally, ensemble lightweight classifiers are trained for classification.
III-B1 Timestep-Wise Feature Bank
In the finetuning step, the parameters of the pretrained DDPM are frozen. The denoising U-Net is able to capture meaningful information from the input data in the reverse process. Leveraging the rich and diverse information captured in the reverse steps is crucial for accurate classification. Therefore, timestep-wise spectral-spatial features are extracted from the intermediate hierarchies of DDPM at different timesteps to construct effective representations that contain diverse information of input HSI. Given an input patch $x_{0}\in\mathcal{R}^{H\times H}$ randomly cropped from HSI and a timestep $t$, Gaussian noise is gradually added to $x_{0}$ according to the variance schedule through diffusion process $q(x_{t}|x_{0})$. The noisy patch $x_{t}$ can be directly obtained by equation 2. Then, the noisy patch $x_{t}$ is fed into the pretrained denoising U-Net to obtain hierarchal features from the U-Net decoder, which implies rich spectral information as well as multi-scale spatial information, as shown in Fig. 2. The features from different layers of the decoder are jointly upsampled to $H\times H$ and then concatenated to get the feature $f_{t}\in\mathcal{R}^{H\times H\times d}$ at timestep $t$. The above process is defined as a function $\mathcal{F}$ and formulated as follows:
$f_{t}=\mathcal{F}(D,x_{t},t)$ (6)
where $D$ is the pretrained denoising U-Net.
For each feature $f_{t_{i}}\in\mathcal{R}^{H\times H\times d}$, we reserve only the vector corresponding to the center pixel indexed as $(\frac{H}{2},\frac{H}{2})$ to obtain the pixel-wise feature $h_{t_{i}}\in\mathcal{R}^{1\times 1\times d}$, which largely reduce computational cost with fewer parameters. Following the above process, we draw $m$ sets of features to construct the timestep-wise feature bank:
=$\mathcal{B}=\{h_{t_{i}}|i\in\{1,...,m\},h_{t_{i}}=center(f_{t_{i}})\}$ (7)
where $t_{1},...,t_{m}$ are sampled from $[0,T]$ at equal intervals. Features extracted at different timestep $t$ tend to capture different information of HSI data since the reverse process is a step-by-step denoising process that tries to predict further information conditioned on the given destructed data at every different timestep.
To further demonstrate this, we visualize the features in 2-dimensional space using t-distributed stochastic neighbor embedding (TSNE) to show the diversity of features at different timesteps $t$. The visualization of the original spectral feature space is shown in Fig. 3(a), and the features in the timestep-wise feature bank at timestep 100, 300, 500, 700, and 900 are visualized in Fig. 3(b)-(f) respectively. As observed, the features in the timestep-wise feature bank show great diversity. Furthermore, with the increase of timestep $t$, the clusters of the features tend to be more loosely structured, and features of the same class tend to be formed into a few large clusters rather than dispersed into multiple smaller clusters. Features at smaller $t$ tend to capture stochastic variation, and features at larger $t$ tend to focus on high-level information, both of which are key to HSI classification with different categories of land covers. Thus the abundant features in the feature bank contain diverse and multi-level information of the input HSI data, exhibiting diverse properties and concerns.
III-B2 Dynamic Feature Fusion Module
For each input HSI patch $x_{0}$, a timestep-wise feature bank is obtained with abundant features containing varied characteristics of the patch. However, the features contain a wealth of information, but not all of it is essential. Using all the features for classification will bring high computational costs and redundant information, and manually selecting timestep-wise features from the feature bank is difficult to characterize spectral-spatial features effectively and robustly. Meanwhile, HSI data have different spectral characteristics for different HSI datasets, resulting in different spectral representation laws of timestep-wise features. Also, data on the same dataset is very different due to its position. Therefore, the dynamic feature fusion module is crucial to be proposed, and is capable of adaptively fusing features of different timesteps from the DDPM to learn effective multi-timestep representations with key information.
To obtain dynamic multi-timestep representations, we design the dynamic feature fusion module to predict the soft dynamic weight of multi-timestep representations. The architecture of the dynamic feature fusion module is illustrated in Fig. 2. Specifically, Let $L=[h_{t_{1}}...h_{t_{m}}]$ denotes the concatenated features from the timestep-wise feature bank as the input of the dynamic weight network. The dynamic weight network consists of two linear layers and outputs the $n$ sets of dynamic weights $w$ with a SoftMax function:
$w$ $=\mathcal{M}(L,W)=SoftMax(a(L,W))$ (8)
$=SoftMax(W_{2}\delta(W_{1}L))$
where $M$ is the dynamic weight network, $W$ is the linear weight of $M$, $\delta$ is ReLU function, $w\in\mathcal{R}^{m\times 1\times 1\times n}$ and $L\in\mathcal{R}^{m\times 1\times 1\times n}$. Inspired by the design of the multi-head attention mechanism in [42], we generate multiple sets of dynamic weights, which allows the representations to focus on different information from different positions and spectral bands. Each dynamic weight determines the proportion of each feature in the corresponding representation. Then, the dynamic weights are used to weight the pixel-wise features by matrix multiplication to obtain $n$ multi-timestep representations:
$r_{i}=w^{T}_{i}L$ (9)
where $r_{i}$ is the i-th representation, $i\in{1,...,n}$. The multi-step representations will be concatenated into the final representation to be classified. Thus, the most appropriate and relevant representation is automatically learned from the feature bank without human intervention, saving time cost and reducing the impact of redundant information. $n$ will be ablated to find the most suitable value balancing the performance and computational cost.
III-B3 Lightweight Ensemble Classifiers
After getting the dynamic representation, a lightweight network is needed to predict the classification label. Inspired by [43], we train an ensemble of lightweight linear classifiers that takes the dynamic pixel representations as inputs and predicts the classification label of each pixel. Specifically, each classifier is trained independently, consisting of two hidden layers with ReLU activation and batch normalization. When testing a sample, the final predicted label is obtained by majority voting of the ensemble of pixel classifiers, as illustrated in Fig. 2. This method brings more stability of prediction with a very small cost since the parameters of each classifier are very limited.
IV Experiments and results
In this section, we first describe the three used well-known HSI datasets, including the Indian Pines dataset, Pavia University dataset and Houston2018 dataset. The experimental setting is then introduced including evaluation metrics, a brief introduction of compared state-of-art methods, and implementation details. Then, we conduct quantitative experiments and ablation analysis to evaluate our proposed method.
IV-A Datasets Description
IV-A1 Indian Pines
The Indian Pines dataset was acquired in 1992 over an area of Indian pines in North-Western Indiana by Airborne Visible Infrared Imaging Spectrometer (AVIRIS) sensor. It consists of $145\times 145$ pixels with a spatial resolution of 20 m and 220 spectral bands in the wavelength range of 400 to 2500 nm. There are 200 bands retained for classification (1-103, 109-149, 164-219) after removing the bands affected by noise. The dataset contains 10249 labeled pixels with 16 categories. We use 10% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table I.
IV-A2 Pavia University
The University of Pavia dataset was collected in 2003 by the reflective optics system imaging spectrometer (ROSIS-3) sensor over a part of the city of Pavia, Italy. The dataset consists of $610\times 340$ pixels with a spatial resolution of 1.3 m and 115 spectral bands in the wavelength range of 430 to 860 nm. 103 out of 115 bands are used for classification after removing 12 noisy bands. The image contains a large number of background pixels and only 42776 labeled pixels divided into 9 classes, including asphalt, meadows, gravel, and so on. We use 5% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table II.
IV-A3 Houston2018
The Houston2018 dataset, identified as 2018 IEEE GRSS DFC dataset, was gathered in 2018 by the National Center for Airborne Laser Mapping (NCALM) over the University of Houston campus and its neighboring urban area, including HSI, multispectral LiDAR, and very high resolution RGB images. The HSI dataset consists of $601\times 2384$ pixels with a spatial resolution of 1 m and 48 spectral bands in the wavelength range of 380 to 1050 nm. It contains 504856 labeled pixels and 20 classes of interest. We use 5% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table III.
TABLE I: Land-cover types, the number of labeled training samples and testing samples of the Indian Pines dataset.
Class Land Cover Type Training Testing
1 Alfalfa 5 41
2 Corn-Notill 143 1285
3 Corn-Mintill 83 747
4 Corn 24 213
5 Grass-Pasture 48 435
6 Grass-Trees 73 657
7 Grass-Pasture-Mowed 3 25
8 Hay-Windrowed 48 430
9 Oats 2 18
10 Soybean-Notill 97 875
11 Soybean-Mintill 245 2210
12 Soybean-Clean 59 534
13 Wheat 20 185
14 Woods 126 1139
15 Buildings-Grass-Trees-Drives 39 347
16 Stone-Steel-Towers 9 84
Total 1024 9225
TABLE II: Land-cover types, the number of labeled training samples and testing samples of the Pavia University dataset.
Class Land Cover Type Training Testing
1 Asphalt 332 6299
2 Meadows 932 17717
3 Gravel 105 1994
4 Trees 153 2911
5 Painted Metal Sheets 67 1278
6 Bare Soil 251 4778
7 Bitumen 67 1263
8 Self-Blocking Bricks 184 3498
9 Shadows 47 900
Total 2138 40638
TABLE III: Land-cover types, the number of labeled training samples and testing samples of the Houston2018 dataset.
Class Land Cover Type Training Testing
1 Healthy Grass 490 9309
2 Stressed Grass 1625 30877
3 Artificial turf 34 650
4 Evergreen trees 680 12915
5 Deciduous trees 251 4770
6 Bare earth 226 4290
7 Water 13 253
8 Residential buildings 1989 37783
9 Non-residential buildings 11187 212565
10 Roads 2293 43573
11 Sidewalks 1702 32327
12 Crosswalks 76 1442
13 Major thoroughfares 2317 44031
14 Highways 493 9372
15 Railways 347 6590
16 Paved parking lots 575 10925
17 Unpaved parking lots 7 139
18 Cars 327 6220
19 Trains 269 5100
20 Stadium seats 341 6483
Total 25242 479614
TABLE IV: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Indian Pines dataset. The best one is shown in bold.
Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
1 65.85 58.54 94.14 63.00 95.12 97.56 72.97 99.10 99.51
2 99.77 76.19 97.84 92.35 97.67 98.05 88.50 96.10 98.89
3 81.66 77.64 97.54 86.86 98.87 98.66 87.20 95.39 99.84
4 96.71 52.11 90.70 88.96 91.55 95.31 84.90 97.24 99.81
5 85.75 93.56 97.75 92.49 96.32 95.17 90.28 94.12 98.85
6 97.87 98.17 99.24 99.12 99.54 99.85 97.97 99.25 100.00
7 100.00 36.00 81.60 52.50 100.00 100.00 56.52 88.46 98.40
8 100.00 98.60 100.00 99.16 100.00 100.00 99.48 100.00 100.00
9 50.00 55.56 74.44 41.18 88.89 100.00 87.50 94.44 100.00
10 35.54 82.86 94.77 93.16 97.71 94.29 86.80 95.84 98.88
11 88.01 90.45 98.87 92.27 98.69 99.37 96.68 99.29 99.63
12 98.13 62.55 97.83 85.44 98.13 96.63 80.83 93.37 98.91
13 99.46 88.65 99.24 99.02 97.28 100.00 100.00 100.00 99.89
14 99.91 99.39 99.18 96.73 99.91 97.89 99.90 95.25 100.00
15 91.35 86.17 93.95 83.41 98.84 97.12 96.80 99.15 99.25
16 86.90 45.24 98.33 93.50 95.54 94.05 84.00 100.00 98.10
OA (%) 87.77 85.42 97.75 92.31 97.47 97.95 92.69 97.02 99.46
AA (%) 86.06 75.10 94.71 84.95 96.57 97.75 88.15 96.00 99.37
$\kappa$ 0.8603 0.8324 0.9743 0.9124 0.9711 0.9766 0.9162 0.9660 0.9938
TABLE V: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$, as well as the accuracies for each class on the Pavia University dataset. The best one is shown in bold.
Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
1 99.68 97.02 98.81 96.21 99.33 99.38 94.20 99.62 100.00
2 99.41 99.97 99.83 99.64 99.92 99.80 99.58 99.98 99.99
3 86.56 92.98 92.45 87.65 98.29 98.35 78.83 91.02 100.00
4 98.18 97.53 98.32 96.64 98.49 99.52 97.53 98.40 99.51
5 99.84 99.06 99.65 99.97 99.53 100.00 100.00 100.00 100.00
6 100.00 99.10 99.43 99.56 100.00 99.75 95.30 98.32 100.00
7 99.84 79.10 99.76 90.30 99.13 99.60 97.42 99.61 100.00
8 100.00 97.34 99.32 94.60 98.05 98.63 96.87 98.36 99.82
9 98.22 95.22 99.82 98.49 95.44 99.33 98.61 99.89 99.96
OA (%) 98.86 97.88 99.10 97.54 99.21 99.53 96.77 99.02 99.94
AA (%) 97.97 95.26 98.60 95.88 98.69 99.37 95.37 98.36 99.92
$\kappa$ 0.9848 0.9719 0.9881 0.9674 0.9915 0.9937 0.9571 0.9870 0.9993
TABLE VI: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Houston2018 dataset. The best one is shown in bold.
Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
1 84.33 82.02 86.30 92.36 79.93 79.50 91.60 88.99 89.41
2 95.98 96.64 95.32 95.08 93.44 96.55 93.92 97.53 96.24
3 96.92 96.00 99.72 96.21 99.66 100.00 97.60 99.23 99.88
4 98.13 95.55 97.49 98.48 96.64 97.62 95.92 97.98 99.26
5 91.26 79.16 86.09 91.87 90.11 95.01 85.79 92.51 96.86
6 99.91 97.51 98.48 99.62 99.62 99.91 98.50 99.58 100.00
7 80.63 71.94 93.68 27.01 85.45 95.65 61.15 96.85 97.63
8 94.29 88.62 91.62 95.42 98.72 99.18 91.15 94.20 99.62
9 99.54 92.80 97.88 98.60 99.09 99.35 95.12 99.07 99.63
10 95.44 71.61 81.72 88.22 91.15 92.64 76.72 88.61 95.89
11 77.86 73.17 69.44 27.01 80.97 85.73 70.22 82.28 93.02
12 15.26 11.10 0.51 31.46 41.69 33.43 4.79 42.76 50.23
13 87.30 69.12 84.59 91.79 94.49 96.79 90.52 93.26 97.72
14 94.97 96.18 89.65 92.91 96.97 99.17 87.23 97.02 99.23
15 95.40 98.98 99.34 99.33 99.30 99.92 99.06 99.59 99.81
16 93.62 90.40 91.00 96.36 97.84 98.15 92.01 96.36 99.69
17 86.33 20.86 0.00 22.78 69.21 90.65 0.00 100.00 94.10
18 96.64 89.05 93.66 91.61 93.07 97.85 90.43 94.24 99.46
19 99.24 95.45 96.92 96.53 97.97 99.84 96.09 99.59 99.89
20 92.77 93.92 99.12 99.77 99.96 100.00 97.27 99.89 99.94
OA (%) 94.84 86.88 91.61 90.65 95.48 96.69 90.34 95.38 98.07
AA (%) 88.79 80.50 82.63 80.75 90.26 92.85 80.76 92.98 95.37
$\kappa$ 0.9325 0.8313 0.8906 0.8784 0.9412 0.9570 0.8751 0.9398 0.9748
IV-B Experimental Setting
IV-B1 Evaluation Metrics
We evaluate the performance of all methods by three widely used indexes: overall accuracy (OA), average accuracy (AA), and Kappa coefficient ($\kappa$).
IV-B2 Comparison with State-of-the-art Methods
To demonstrate the effectiveness of our proposed method, we compare our classification performance with several state-of-the-art approaches, including CNN-based and transformer-based supervised methods, and unsupervised methods. We use the most effective setting for each of these methods.
  • The 2-D CNN architecture contains three 2-D convolution blocks and a softmax layer. Each convolution block consists of a 2-D convolution layer, a BN layer, an avg-pooling layer, and a ReLU activation function.
  • The 3-D CNN contains three 3-D convolution blocks and a softmax layer. Each 3-D convolution block consists of a 3-D convolution layer, a BN layer, a ReLU activation function, and a 3-D convolution layer with step size 2.
  • The SSRN[14] is a spectral-spatial residual network based on 3-D CNN and residual connection. Spatial residual blocks and spatial residual blocks are designed to extract discriminative features from HSI data.
  • For SF[16], there are four encoder blocks in the ViT-based network. In the transformer framework of SF, group-wise spectral embedding and cross-layer adaptive fusion modules are adopted to capture local spectral representations from neighboring bands.
  • The SSFTT[17] systematically combines CNN network and transformer structure to exploit spectral-spatial information in the HSI, with a Gaussian weighted feature tokenizer module making the samples more separable.
  • The GAHT[18] is a end-to-end group-aware transformer method with three-stage hierarchical framework.
  • The 3DCAE[21] is an unsupervised method using the encoder-decoder backbone with 3D convolution operation to learn spectral-spatial features.
  • The UMSDFL[22] is an unsupervised method using encoder and decoder with convolutional layers to learn spectral-spatial features. A clustering branch and a multi-layer fusion module is designed to enhance the features.
IV-B3 Implementation Details
The proposed Diff-HSI was implemented using the Pytorch framework. The patch size is set to $48\times 48$, and the dimension of PCA is set to 10. In the diffusion-pretraining procedure, we use Kullback-Leibler Divergence Loss as the loss function. And the Adam optimizer is adopted with a batch size of 128 and a learning rate of 1e-4. In the second stage, the cross-entropy loss is used in lightweight classifiers. $m$, $n$, and $k$ are set to be 20, 3, and 5, respectively. We adopt the Adam Optimizer and the Cosine Annealing as our training schedule. The original learning rate and minimum learning rate are set to be 1e-4 and 5e-6, respectively. The number of epochs is set to 100 for all datasets. Fairly, we calculate the results by averaging the results of ten repeated experiments with different training sample selections.
IV-C Quantitative Results and Analysis
Quantitative classification results in terms of class-specific accuracy, OA, AA, and $\kappa$ of the compared methods on the Indian Pines, PaviaU, and Houston 2018 datasets are listed in Table IV, V, and VI, respectively. And the classification maps of all methods are shown in Fig. 4, 5 and 6.
Compared with other methods, our proposed Diff-HSI achieves the highest OA, AA, and $\kappa$ on the three datasets. According to the results, the CNN-based supervised methods obtain good performance owing to the ability of capturing local spatial information. Besides, since transformers are capable of capturing sequential information, transformer-based supervised methods also achieve competitive performance. The GAHT method achieves the best results in CNN-based and transformer-based supervised methods because it balances the local information and the global dependencies. However, the methods that learn explicit mappings have bottlenecks in learning label-agnostic information and complex relations, which limits performance. Some unsupervised methods are proposed to tackle the problem by learning representative features through an encoder-decoder network. Limited by the model architecture, the features learned are not discriminative and rich enough for HSI classification lacking high-level information. Therefore, the performance of these methods is even lower than that of some explicit learning methods. Our proposed Diff-HSI introduces the diffusion model to HSI classification, which has the powerful ability to model complex relations. Furthermore, the features learned by our method are of great diversity, enriching the representation capability. Thus, our Diff-HSI outperforms all the previous methods on the three datasets. Notably, the classification performance of the Houston2018 Dataset is largely improved compared with the previous SOTA method in terms of OA (98.07% versus 96.69%), AA (95.37% versus 92.98%), and $\kappa$ (0.9748 versus 0.9570), which especially demonstrate our effectiveness.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
Figure 4: Classification maps obtained by different methods on the Indian Pines dataset. (a) Ground truth. (b) 2-D CNN (OA=87.77%). (c) 3-D CNN (OA=85.42%). (d) SSRN (OA=97.75%). (e) SF (OA=92.31%). (f) SSFTT (OA=97.47%). (g) GAHT (OA=97.95%). (h) 3DCAE (OA=92.69%). (i) UMSDFL (OA=97.02%). (j) Diff-HSI (OA=99.46%).
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
Figure 5: Classification maps obtained by different methods on the Pavia University dataset. (a) Ground truth. (b) 2-D CNN (OA=98.86%). (c) 3-D CNN (OA=97.88%). (d) SSRN (OA=99.10%). (e) SF (OA=97.54%). (f) SSFTT (OA=99.21%). (g) GAHT (OA=99.53%). (h) 3DCAE (OA=96.77%). (i) UMSDFL (OA=99.02%). (j) Diff-HSI (OA=99.94%).
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
Figure 6: Classification maps obtained by different methods on the Houston2018 dataset. (a) Ground truth. (b) 2-D CNN (OA=94.84%). (c) 3-D CNN (OA=86.88%). (d) SSRN (OA=91.61%). (e) SF (OA=90.65%). (f) SSFTT (OA=95.48%). (g) GAHT (OA=96.69%). (h) 3DCAE (OA=90.34%). (i) UMSDFL (OA=95.38%). (j) Diff-HSI (OA=98.07%).
IV-D Ablation Studies
In this section, we analyze the effect of the components in our method.
IV-D1 Ablation for Set Number of Dynamic Weights
Our dynamic feature fusion module contains $n$ sets of dynamic weights which impact the classification performance. The change of results by changing $n$ is shown in Tabel VII. We reveal that the best performance is achieved when $n$ is taken as 3. With smaller $n$, only part of the useful information is obtained from the feature bank, which is insufficient for classification. On the other hand, larger $n$ will bring more parameters and excess information.
IV-D2 Ablation for Pretraining Steps
We validate the effectiveness of pretraining on the Indian Pines dataset in terms of OA, AA, and $\kappa$. As shown in Table VIII, only 10k steps pretraining brings dramatic improvement (more than 30% OA) to the final classification performance. Furthermore, as the pretraining steps increase, the performance continues to rise to the best at around 40k steps.
IV-D3 Sensitivity Analysis of Timestep T
To analyze the features extracted from different timestep $t$, we record the change of the classification performance when changing $t$. For easy understanding, we choose 4 of the 16 classes to show their changes, as illustrated in Fig. 7. The performance for each class behaves differently as the $t$ increases. Features of larger $t$ are more sensitive to class ”corn-notill” and features of smaller $t$ are more informative to class ”woods”. For class ”Grass-pasture-mowed” and class ”corn”, features extracted at the intermediate $t$ is the most discriminative. Thus, an appropriate set of $t$ is vital for accurate performance.
TABLE VII: Ablation analysis of the proposed dynamic feature fusion module with a different number of dynamic weight sets in terms of OA, AA, and $\kappa$ on the Indian Pines dataset.
Metric The number of the sets of dynamic weights $n$
1 2 3 4
OA (%) 99.36 99.38 99.46 99.43
AA (%) 99.20 99.20 99.37 99.28
$\kappa$ 0.9918 0.9930 0.9938 0.9935
TABLE VIII: Effect of Pretraining Steps in terms of OA, AA, and $\kappa$ on the Indian Pines dataset.
Metric The number of pretraining steps
0 10k 20k 30k 40k 50k
OA (%) 68.04 97.53 98.40 99.39 99.46 99.45
AA (%) 48.95 96.64 97.76 99.30 99.37 99.39
$\kappa$ 58.51 0.9718 0.9817 0.9930 0.9938 0.9936
Figure 7: Variation of classification results for four classes with the change of timestep on the Indian Pines dataset.
TABLE IX: Classification performance with different sets of timestep $t$ on the Indian Pines dataset.
Timestep Sets $X$ Metric
OA (%) AA (%) $\kappa$
$[100,200,300]$ 98.81 98.72 0.9874
$[700,800,900]$ 98.77 98.68 0.9872
$[100,500,900]$ 98.87 98.83 0.9886
Dynamic Feature Fusion 99.46 99.37 0.9938
Table IX shows the results with several sets of manually assigned $t$. Specifically, given a timestep set $X=[t_{1},t_{2},t_{3}]$, we extract the features at timestep $t_{1}$, $t_{2}$, and $t_{3}$ to construct the pixel-wise representation for classification. As Table IX presents, the performance is sensitive to the choice of $t$. $X$ with a large $t$, a small $t$ and a medial $t$ performs better than $X$ with only smaller $t$ or larger $t$. Because the former contains a wider range of $t$ leading to more diverse features which make the representation more general for samples of all classes. However, representation obtained by the manually designed method is not always the most appropriate for each sample, since it sacrifices some specificity to improve generalization. So we propose the dynamic feature fusion module to customize $T$ for each sample which makes it more accurate for classification. The results in Table IX demonstrate that our dynamic feature fusion module learns more effective representation compared with the manually selected method.
Figure 8: Classification results of different patch sizes on the Indian Pines, Pavia University, and Houston2018 Dataset.
Figure 9: Classification results of different PCA components on the Indian Pines, Pavia University, and Houston2018 Dataset.
IV-E Parameter Analysis
In this section, we analyze the effect of various parameters that influence classification performance by training our proposed Diff-HSI in the same experimental setting as Section IV-B with different parameters.
IV-E1 Effect of Different Patch Size
First, we discuss the effect of different patch sizes on classification performance by two indexes, OA and AA. As shown in Fig. 8, the patch size varies from 24$\times$24 to 60$\times$60. as the patch size increases, the performance first increases and then decreases. The best performance is obtained when the patch size is 48$\times$48 with OA of 99.39% and AA of 99.30%. Too small patches contain insufficient spatial information and too large patches reduce the attention to detailed structures. Thus, we choose 48$\times$48 to be the patch size for the proposed Diff-HSI.
IV-E2 Effect of Different PCA Components
This section analyzes the influence of the number of PCA components, which determines how much spectral information is retained in the compressed data. As the number of PCA components increases, more spectral information is retained while more computational cost and more redundant information are brought. Since each dataset has a different number of channels, the range of PCA components is different for the three datasets. Assuming that $N$ is the channel number of a dataset, the number of PCA components varies from $N/20$ to $N/5$. According to the results shown in Fig. 9, the best performance is achieved at PCA components of $N/8$.
V Conclusion
HSI contains rich spectral-spatial information and complex relations, which are critical for classification tasks. We propose the Diff-HSI method to learn discriminative spectral-spatial features in an unsupervised manner using the diffusion model. Most of the existing works perform HSI classification in an explicit way based on CNN or Transformer. These methods often heavily rely on the number of labeled data and are insufficient to model complex relations, limiting the improvement in performance. Some works explore unsupervised learning for HSI classification through an encoder-decoder feature learning method. However, limited by the feature-extracting capability of the model, they mainly tend to capture low-level features, lacking high-level features which are significant for discriminative spectral-spatial feature representation. Our proposed method leverages the powerful implicit learning ability of the diffusion model to learn effective spectral-spatial features with high degrees of freedom, constructing the step-wise feature bank. Furthermore, we design a dynamic feature fusion model to obtain multi-timestep representation from the feature bank automatically. Quantitative experiments on three HSI datasets demonstrate that our proposed Diff-HSI outperforms state-of-art supervised and unsupervised methods.

### v2

@@
-When Hyperspectral Image Classification Meets Diffusion Models: An Unsupervised Feature Learning Framework
+Exploring Multi-Timestep Multi-Stage Diffusion Features for Hyperspectral Image Classification
 Abstract
-Learning effective spectral-spatial features is important for the hyperspectral image (HSI) classification task, but the majority of existing HSI classification methods still suffer from modeling complex spectral-spatial relations and characterizing low-level details and high-level semantics comprehensively.
-As a new class of record-breaking generative models, diffusion models are capable of modeling complex relations for understanding inputs well as learning both high-level and low-level visual features.
-Meanwhile, diffusion models can capture more abundant features by taking advantage of the extra and unique dimension of timestep $t$.
-In view of these, we propose an unsupervised spectral-spatial feature learning framework based on the diffusion model for HSI classification for the first time, named Diff-HSI.
-Specifically, we first pretrain the diffusion model with unlabeled HSI patches for unsupervised feature learning, and then exploit intermediate hierarchical features from different timesteps for classification.
-For better using the abundant timestep-wise features, we design a timestep-wise feature bank and a dynamic feature fusion module to construct timestep-wise features, adaptively learning informative multi-timestep representations.
-Finally, an ensemble of linear classifiers is applied to perform HSI classification.
-Extensive experiments are conducted on three public HSI datasets, and our results demonstrate that Diff-HSI outperforms state-of-the-art supervised and unsupervised methods for HSI classification.
+The effectiveness of spectral-spatial feature learning is crucial for the hyperspectral image (HSI) classification task.
+Diffusion models, as a new class of groundbreaking generative models, have the ability to learn both contextual semantics and textual details from the distinct timestep dimension, enabling the modeling of complex spectral-spatial relations in HSIs.
+However, existing diffusion-based HSI classification methods only utilize manually selected single-timestep single-stage features, limiting the full exploration and exploitation of rich contextual semantics and textual information hidden in the diffusion model.
+To address this issue, we propose a novel diffusion-based feature learning framework that explores Multi-Timestep Multi-Stage Diffusion features for HSI classification for the first time, called MTMSD.
+Specifically, the diffusion model is first pretrained with unlabeled HSI patches to mine the connotation of unlabeled data, and then is used to extract the multi-timestep multi-stage diffusion features.
+To effectively and efficiently leverage multi-timestep multi-stage features, two strategies are further developed.
+One strategy is class & timestep-oriented multi-stage feature purification module with the inter-class and inter-timestep prior for reducing the redundancy of multi-stage features and alleviating memory constraints.
+The other one is selective timestep feature fusion module with the guidance of global features to adaptively select different timestep features for integrating texture and semantics.
+Both strategies facilitate the generality and adaptability of the MTMSD framework for diverse patterns of different HSI data.
+Extensive experiments are conducted on four public HSI datasets, and the results demonstrate that our method outperforms state-of-the-art methods for HSI classification, especially on the challenging Houston 2018 dataset.
+The codes are available at https://github.com/zjyaccount/MTMSD.
 Index Terms:
-Hyperspectral image classification, denoising diffusion probabilistic model, unsupervised feature learning.
+Hyperspectral image classification, denoising diffusion probabilistic model, multi-timestep multi-stage features, feature purification, feature selection.
 I Introduction
-HYPERSPECTRAL image (HSI) classification plays a crucial role in remote sensing [1, 2, 3, 4], as it aims to distinguish each pixel’s category in hyperspectral data by using dense and detailed electromagnetic spectral information [5].
-The main objective is to effectively discriminate objects of interest by capturing subtle variations in the spectral signatures associated with their pixels [6].
-To this end, a large number of methods are developed, relying on hand-crafted features [7, 8, 9, 10, 11].
-However, these works are designed to extract features manually, leading to limited discriminative information extraction from the high spectral variability [12].
-The rapid advancement of deep learning has led to the emergence of supervised approaches, a typical category of HSI classification methods, specifically utilizing deep convolutional neural networks (CNNs) [13, 14, 15] and transformers [16, 17, 18, 19].
-Although these previous supervised methods have already achieved promising results in HSI classification, they still have limitations in spectral-spatial feature extraction.
-Firstly, previous supervised methods train the CNN or transformer network to learn explicit mappings which leads to poor performance to model complex spectral-spatial relations.
-Secondly, supervised methods tend to prioritize features from labeled areas while HSIs have large unlabeled regions that contain label-agnostic information.
-They result in a preference for learning labeling priors and limited general representation generation that characterize spectral-spatial relations.
-Additionally, the limited availability of labeled samples in HSI datasets poses constraints on previous supervised methods.
-((a))
-((b))
-Figure 1: Different unsupervised feature learning frameworks including (a) existing unsupervised feature learning and (b) ours proposed Diff-HSI that extract timestep-wise features and construct multi-timestep representations by dynamic fusion for the first time.
-As another category of HSI classification methods, [20, 21, 22, 23] is designed to learn spectral-spatial features from HSIs in an unsupervised manner.
-Typically, these methods leverage an encoder-decoder network trained in an unsupervised manner for HSI reconstruction and extract the spectral-spatial feature from the intermediate feature maps, depicted in Fig.
-1(a).
-Although these methods mine implicit spectral attributes and spatial relations through unsupervised feature learning, the reconstruction networks they designed mainly capture low-level reconstruction features such as texture and edge, lacking high-level features which are necessary for discriminative representations to model the spectral-spatial relations comprehensively [24].
-Learning rich representations and modeling the complex spectral-spatial relations from high dimensions are significant and should be considered for HSI classification.
-Most recently, diffusion models [25, 26, 27] have emerged as new generative models with superior performance in generation and reconstruction tasks, and have started to be explored in other computer vision tasks like semantic segmentation [28, 29, 30].
-Diffusion models have the ability in the reconstruction task, so they are well-suited as a framework for unsupervised feature learning.
-Furthermore, there exist two appealing properties of using diffusion models [31, 32, 33] for unsupervised feature learning of hyperspectral data: 1) diffusion models learn implicit mappings, which have superior representation ability on complex problems via simple expressions and extract both high-level and low-level features from images; 2) the stepwise reverse denoising process of diffusion models are modeled as an iterative optimization process optimized by Langevin dynamics.
-This process brings a large number of degrees of freedom and tries to predict further information conditioned on the given noise-destructed data at every timestep, leading to better generalization to complex problems.
-When applying diffusion models to unsupervised feature learning of HSI data, it not only learns both high-level and low-level features but also extracts spectral-spatial features in the dimension of timestep $t$, as shown in Fig.
-1(b).
-Naturally, it is crucial to explore how to exploit abundant timestep-wise features more effectively and efficiently.
-On the one hand, the timestep-wise features are diverse and informative and may focus on different information.
-The shallow timestep-wise features are more informative for fine details, while deeper ones are more concerned about high-level semantics and global information [25, 34].
-On the other hand, HSIs from various datasets acquired by different sensors exhibit distinct spectral characteristics, leading to variations in the spectral representation patterns of timestep-wise features extracted from different HSI datasets.
-Meanwhile, the abundant timestep-wise features pose a challenge in terms of computation overheads.
-In addition, manually selecting timestep-wise features to characterize spectral-spatial features effectively is struggle.
-Therefore, a dynamic feature fusion module that adaptively fuses features of different timesteps for different HSI datasets is desired.
-In view of these, we propose the unsupervised spectral-spatial feature learning framework based on the diffusion model for HSI classification, named Diff-HSI.
-More specifically, the proposed framework first pretrain a diffusion model with unlabeled HSI patches for unsupervised feature learning, which is able to automatically learn label-agnostic knowledge and mine the connotation of unlabeled data that reveals the complex spectral-spatial dependencies compared to conventional supervised methods.
-Then, we extract features from different hierarchies from the pretrained denoising U-Net decoder to construct the timestep-wise feature bank, which contains various informative features from different timestep $t$.
-To effectively harness the features from the timestep-wise feature bank and softly learn the proper timestep-wise feature combination for different HSI datasets, we propose a dynamic feature fusion module.
-This module is designed to adaptively fuse hierarchical features from different timesteps, generating multi-timestep representations enriched with comprehensive spectral-spatial information.
-Ultimately, an ensemble of linear classifiers is employed to leverage these dynamic representations for accurate HSI classification.
+HYPERSPECTRAL image (HSI) classification plays a crucial role in remote sensing, as it aims to distinguish each pixel’s category in hyperspectral data by using dense and detailed electromagnetic spectral information [1, 2].
+HSI classification has a wide range of applications in environmental monitoring [3], resource management [4], agriculture disaster response [5], military defense [6], etc.
+The extraction of spectral-spatial features from HSIs plays a significant role in HSI classification tasks.
+Initially, machine learning-based feature extraction methods have been developed [7, 8, 9, 10, 11].
+However, these approaches rely on manual feature engineering, resulting in limited extraction of discriminative information from the highly variable spectral data [12].
+In recent years, the rapid advancement of deep learning has paved the way for neural network-based approaches for feature extraction, specifically leveraging convolutional networks (CNNs) [13, 14, 15] and transformer models [16, 17, 18, 19].
+These neural network-based methods excel at automatically learning valuable spectral-spatial features from labeled HSI data, thus achieving promising results in HSI classification.
+Further, unsupervised methods such as [20, 21, 22] have been designed to excavate spectral-spatial features from unlabeled HSIs.
+Typically, these methods employ an encoder-decoder network trained in an unsupervised manner for HSI reconstruction, thus extracting spectral-spatial features from unlabeled HSI data.
+Leveraging unsupervised feature learning enables deep mining of large amounts of unlabeled regions in HSI, which contain a wealth of label-agnostic information, thereby promoting the HSI classification task.
+Figure 1: Overview of existing diffusion-based feature learning frameworks for the HSI classification task.
+Our method can fully explore and exploit rich contextual semantics and textual features hidden in the diffusion model.
+Recently, diffusion models [23, 24, 25] have emerged as powerful models with superior performance in generation and reconstruction tasks.
+These models have also been explored in various computer vision tasks such as semantic segmentation [26, 27, 28].
+Distinguished from traditional deep neural networks, diffusion models employ a stepwise reverse denoising process, formulated as an iterative optimization procedure optimized by Langevin dynamics [29].
+This approach introduces a multitude of degrees of freedom for feature learning and predicts additional information conditioned on the given noise-corrupted data at each timestep.
+Consequently, diffusion models can implicitly capture both high-level and low-level visual concepts, facilitating better generalization and modeling of complex spectral-spatial relations [30, 31].
+Therefore, one recent work has proposed using diffusion models for HSI classification [32], which employs unsupervised feature learning to extract diffusion features.
+However, the diffusion features used in [32] are extracted solely from a single timestep and a single stage of the denoising U-Net, and these selections about the timestep and stage are manually determined based on extensive experimentation with each dataset, as shown in Fig.
+1.
+Firstly, relying on single-timestep features from a single U-Net layer inevitably results in the loss of abundant spectral-spatial information and limits the effectiveness of modeling spectral-spatial relations.
+Secondly, the manual selection approach lacks generality and adaptability to diverse HSI datasets exhibiting specific spectral representation patterns.
+Naturally, it is crucial to explore how to exploit abundant multi-timestep multi-stage features extracted from the whole stage of denoising U-Net more effectively and efficiently.
+First, multi-timestep features extracted from diffusion models are diverse and focus on different information.
+The shallow timestep-wise features are more informative for textual details, while deeper ones are more concerned about high-level semantics and global information [23, 33].
+For modeling complex spectral-spatial relations, textual features depict the spatial distribution and changing patterns of ground objects, and contextual semantics represent the spectral attributes and the content of ground objects.
+Thus, leveraging multi-timestep features can integrate both contextual semantics and textual features to build comprehensive spectral-spatial representation for better performance.
+Second, multi-stage features, as compared to single-stage features, encapsulate a rich hierarchy of information that contains richer semantic and reconstruction characteristics, facilitating the modeling of spectral-spatial relations.
+Moreover, [32] also demonstrates that various stage features improve classification performance with different degrees across various datasets.
+Although multi-timestep multi-stage features are desired, HSIs from various datasets acquired by different sensors exhibit distinct spectral-spatial characteristics, leading to variations in the spectral-spatial representations.
+Moreover, for different regions in the same HSI, the emphasis on texture and semantics varies, leading to distinct preferences in the selection of timestep $t$.
+Meanwhile, multi-stage features are numerous and comprise both redundant semantics and reconstruction information along the channel dimension, posing a challenge in terms of memory constraints when training and inferring on large datasets.
+In view of these, we propose a novel diffusion-based feature learning framework that explores Multi-Timestep Multi-Stage Diffusion features for HSI classification for the first time, named MTMSD.
+More specifically, the proposed framework first pretrain a diffusion model with unlabeled HSI patches for diffusion feature learning, mining the connotation of unlabeled data that reveals the complex spectral-spatial dependencies.
+Then, we extract multi-timestep multi-stage diffusion features from the pretrained denoising U-Net decoder and construct the timestep-wise center and global feature bank by center extraction and average pooling.
+After that, two strategies are further developed in MTMSD to leverage multi-timestep multi-stage diffusion features effectively and efficiently.
+First, to reduce the redundancy of multi-stage features and maintain efficiency, we propose to perform class & timestep-oriented multi-stage feature purification on multi-stage features in the timestep-wise center feature bank with the inter-class and inter-timestep prior.
+Second, to effectively harness multi-timestep features and softly learn the proper timestep-wise feature combination for different datasets, we propose the selective timestep feature fusion module.
+This module is designed to adaptively select different timestep center features with the guidance of related global features, and fuse them for multi-timestep multi-stage selective representations that integrate contextual semantics and textual features to model comprehensive spectral-spatial relations.
+Ultimately, an ensemble of linear classifiers is employed for accurate HSI classification.
 To summarize, our contributions are listed as follows.
   • 1)
-    As a pioneer work, we introduce diffusion models to HSI classification and propose a diffusion-based unsupervised feature learning framework called Diff-HSI, which can implicitly learn both high-level and low-level features to model complex spectral-spatial relations.
+    For modeling complex spectral-spatial relations, we propose a novel diffusion-based framework that explores multi-timestep multi-stage diffusion features for HSI classification.
+To the best of our knowledge, this is the first work to learn and exploit multi-timestep multi-stage diffusion features for diffusion-based HSI classification.
   • 2)
-    To leverage abundant timestep-wise features effectively and efficiently, we construct a timestep-wise feature bank with hierarchical features from different timesteps, and design a dynamic feature fusion module to fuse these timestep-wise features adaptively.
-The above process learns the soft timestep-wise feature combination for each labeled patch of different HSI datasets and constructs comprehensive multi-timestep representations.
+    We design the novel class & timestep-oriented multi-stage feature purification module.
+It adaptively selects significant channels of multi-stage diffusion features from both inter-class and inter-timestep aspects to reduce redundant information and maintain computational efficiency.
   • 3)
-    Compared with several state-of-the-art supervised and unsupervised methods, experimental results demonstrate that our proposed method achieves superior classification accuracy on three public HSI datasets.
+    We propose to perform selective timestep feature fusion on multi-timestep diffusion features.
+This module allows each labeled patch of different datasets to adaptively select different timestep center features with the guidance of related global features, and fuse them for comprehensive multi-timestep multi-stage selective representations that integrate contextual semantics and textual information.
+  • 4)
+    Compared with several state-of-the-art HSI classification methods, experimental results demonstrate that our proposed method achieves significant classification accuracy on four public HSI datasets, especially on the challenging Houston 2018 dataset.
 The remainder of this paper is organized as follows.
 Section II describes related work.
-In Section III, our proposed Diff-HSI is introduced in detail.
-Section IV conducts extensive experiments on three HSI datasets to demonstrate the effectiveness of the proposed method.
+In Section III, our proposed MTMSD is introduced in detail.
+Section IV conducts extensive experiments on four HSI datasets to demonstrate the effectiveness of the proposed method.
 Finally, some conclusions are drawn in Section V.
 II Related Work
-II-A Supervised-based HSI Classification Methods
-Due to the remarkable breakthroughs achieved by deep learning in various computer vision tasks, many progressive deep learning-based networks have been widely utilized for supervised HSI classification methods, e.g., autoencoders (AEs) [21], deep belief networks (DBNs) [35] and recurrent neural networks (RNNs) [36].
-Among these, CNNs draw significant attention with their excellent feature extraction capability and become mainstream in HSI classification [13, 14, 15].
-CNNs can effectively extract spectral-spatial features through their ability to extract spatially structural information and locally contextual information.
-As the vision transformer rises in the field of computer vision, researchers begin to explore the value of the transformer in HSI classification and find it outperforms CNN-based methods in solving the problem of long-term spectral information dependencies [16, 17, 18, 19].
-Although these supervised methods have achieved promising performance, they still require large amounts of labeled HSI samples, and are limited to modeling complex relations and capturing label-agnostic information.
-II-B Unsupervised-based HSI Classification Methods
-Unsupervised feature learning is a feature extraction paradigm to automatically learn feature representations from the input data without any annotated information.
-Leveraging unsupervised feature learning on unlabeled HSI patches is a solution to the limited labeled samples of HSI datasets.
-Typically, the commonly used unsupervised feature learning methods in HSI classification are based on the encoder–decoder paradigm, where an autoencoder-like network encodes the input HSI patches into a refined feature and then reconstructs the feature to initial HSI data by a decoder.
+II-A HSI Classification
+HSI classification is an important research topic in the area of remote sensing.
+Since HSI classification aims to distinguish each pixel’s category in hyperspectral data using dense electromagnetic spectral information [1], a large number of handcrafted feature-based methods have been designed for HSI classification [7, 8, 9, 10, 11].
+Several works adopt morphological profiles (MPs) for manually extracting spectral-spatial features from HSIs.
+They achieve good classification results using MPs as input vectors with a support vector machine classifier.
+Subspace-based learning, such as sparse representation and manifold learning, is another common feature extraction strategy for HSI classification.
+These methods transform the high-dimensional original space using a low-dimensional subspace representation to learn spectral-spatial information.
+Due to the remarkable breakthroughs achieved by deep learning in various computer vision tasks, many progressive deep learning-based networks have been widely utilized for HSI classification methods.
+Among these, CNNs draw significant attention with their feature extraction capability to extract spatially structural information and locally contextual information and become mainstream in HSI classification [13, 14, 15].
+Based on the spectral and spatial attention modules, Zhu et al.
+[14] embed a residual block into a sequential spectral-spatial feature learning network.
+This architecture not only mitigates the risk of overfitting but also enhances classification performance.
+However, CNNs have the challenge of modeling long-term spectral information dependencies.
+To address this defect, researchers explore the value of the transformer and widely leverage it in HSI classification [16, 17, 18, 19].
+Hong et al.
+[16] consider the spectral sequence of neighboring bands and design a pure transformer-based SpectralFormer (SF) backbone network, representing sequence attributes of spectral signatures.
+Sun et al.
+[17] extract high-level semantic features by introducing a Gaussian weighted token module into the transformer architecture, achieving promising performance in both classification accuracy and computational complexity.
+In addition to supervised feature learning, unsupervised feature learning aims to learn feature representations from the input data without any annotated information, providing a solution to the limited labeled samples of HSI datasets.
+Typically, the commonly used unsupervised feature learning methods in HSI classification are based on the encoder–decoder paradigm, where an autoencoder-like network encodes the input HSI patches into a purified feature and then reconstructs the feature to initial HSI data by a decoder network.
 Mou et al.
@@
 [21] design a multi-scale CNN-based unsupervised feature learning framework, with two branches of decoder and clustering optimized by the error feedback of image reconstruction and pseudo-label classification.
-In the aforementioned approaches, the feature representations learned from the unlabeled HSI patches contain more spectral-spatial information than supervised methods; however, they mainly focus on low-level visual concepts and ignore high-level feature extraction for HSI classification, which is solved in our methods.
-II-C Diffusion Models
+More recently, with the rise of diffusion models, one recent work propose a diffusion-based HSI classification method.
+Chen et al.
+propose SpectralDiff [32] that extracts the spectral-spatial diffusion features from spectral–spatial denoising network and directly feeds them into the attention-based classification network for classification.
+However, they only use a single timestep feature from a single stage of the denoising U-Net, which is manually selected according to extensive experiments on each dataset.
+This results in a lack of information to model spectral-spatial relations and poor robustness to diverse HSI datasets with diverse spectral characteristics.
+In our work, we propose a novel diffusion-based HSI classification framework that explores multi-timestep multi-stage diffusion features.
+Through multi-stage feature purification and selective timestep fusion, our MTMSD enables adaptive integration of both contextual semantics and textual features to model complex spectral-spatial relations and generality for diverse patterns of different HSI data.
+II-B Diffusion Models
 Diffusion models are a class of probabilistic generative models that progressively inject a standard Gaussian noise, then learn a model to reverse this process for sample generation [23, 24, 25].
 Current research on diffusion models is mostly based on three formulations: denoising diffusion probabilistic models (DDPMs) [23], score-based generative models [31], and stochastic differential equations [35].
-DDPMs are the mainstream diffusion models, and a large number of recent works based on DDPMs have made DDPMs increasingly powerful in terms of generative quality and diversity over other generative models [38].
+Among them, DDPMs are the mainstream diffusion models, and a large number of recent works based on DDPMs have made DDPMs increasingly powerful in terms of generative quality and diversity over other generative models [36].
 Meanwhile, DDPMs have been widely used in several applications, including super-resolution [37], inpainting [38], and point cloud generation [39].
-However, to the best of our knowledge, few works successfully adopt them to HSI classification.
-When applying diffusion models to HSI classification, it can not only implicitly learn both high-level and low-level features but also extract spectral-spatial features in the dimension of timestep $t$, leading to improvement in modeling complex spectral-spatial relations.
-Figure 2: Overview of our proposed Diff-HSI.
+Recently, [40] proposes a simple diffusion-based semantic segmentation approach that exploits 3-timestep multi-stage features with manually selected timesteps and also proves that diffusion features capture high-level semantic information for semantic segmentation.
+However, a fixed timestep set results in suboptimal and non-generic features for different datasets.
+Additionally, such a simple utilization of diffusion features for segmentation may lead to inefficiencies due to redundant information and ineffectiveness in building semantic representations from diffusion features.
+Differently, we focus on the HSI classification task, and dynamically select the timesteps and stages of multi-timestep multi-stage features ranging from low to high, integrating textural features and semantics to build comprehensive spectral-spatial representations.
+II-C Deep Feature Selection
+Feature selection [41], an essential process in deep-learning-based computer vision, plays a pivotal role in improving model performance.
+The goal of feature selection is to retain informative and refined features from the original features, thereby reducing dimensionality and computational complexity.
+For different images, the important features vary, and feature selection assists models in adapting to data from diverse domains, extracting significant and refined features.
+Early feature selection methods select features with input-dependent soft attention.
+Hu et al.
+[42] introduce a channel attention module named SENet to adaptively recalibrate channel-wise features by exploiting the inter-channel relations.
+Similar to channel attention, spatial attention methods GENet [43] is designed to enhance a network’s capacity for context information modeling via spatial masks.
+Building on these, Woo et al.
+[44] combine both channel and spatial attention, introducing the CBAM method.
+Additionally, feature selection on different kernel features is also a self-adaptive and effective mechanism.
+Li et al.
+[45] propose SKNet to select the features extracted from different convolutional kernels using softmax attention along the channel dimension.
+Different from the above methods, our framework is the first to use feature selection on multi-timestep multi-stage features for HSI classification.
+In detail, MTMSD creatively employs inter-class and inter-timestep priors for multi-stage feature purification in the channel dimension, and concurrently conducts feature selection on multi-timestep features with the guidance of global information.
+Figure 2: Overview of our proposed MTMSD.
 The method consists of two steps.
-Step 1: We pretrain the DDPM with HSI patches in an unsupervised manner and then freeze the parameters.
-Step 2: We construct a timestep-wise feature bank by extracting hierarchical features from the fixed pretrained U-Net decoder in terms of different timestep $t$.
-In order to softly fuse the features, a dynamic feature fusion module is designed to obtain effective multi-timestep representations adaptively.
-Finally, an ensemble of lightweight classifiers is trained by leveraging the multi-timestep representations of label-limited HSI patches to perform HSI classification.
+Step 1: We pretrain the DDPM with HSI patches in an unsupervised manner for diffusion feature learning.
+Step 2: We extract multi-timestep multi-stage diffusion features from the pretrained denoising U-Net decoder and construct the timestep-wise center and global feature bank by center extraction and average pooling.
+To effectively and efficiently leverage multi-timestep multi-stage features, we first perform class & timestep-oriented multi-stage purification on multi-stage features in the timestep-wise center feature bank, and then, we perform selective timestep feature fusion with global-feature guidance on the purified timestep-wise center feature bank.
+Classification is performed through an ensemble of lightweight classifiers.
 III Method
-Our proposed Diff-HSI aims to learn effective spectral-spatial representations in an unsupervised manner via denoising diffusion probabilistic model (DDPM) for HSI classification.
+Our proposed MTMSD is a novel diffusion-based feature learning framework and aims to explore multi-timestep multi-stage diffusion features effectively and efficiently for modeling spectral-spatial relations comprehensively.
 The framework is shown in Fig.
 2.
-Our method consists of two key steps.
-The first step is to pretrain the DDPM with unlabeled patches to learn complex spectral-spatial distributions and dependencies of HSI data.
-After pretraining, we freeze the parameters of the pretrained DDPM.
-The second step of our Diff-HSI is a supervised learning process aiming to fully extract and dynamically fuse the timestep-wise features from the pretrained denoising U-Net for classification.
-In order to capture the characteristics of different objects in HSI data, we first extract abundant spectral-spatial features from different timestep $t$ to construct the timestep-wise feature bank.
-Then, we design a dynamic feature fusion module that fuses the features in the feature bank dynamically and generates multi-timestep representations with discriminative information for classification.
-Finally, the representations are fed into an ensemble of lightweight classifiers to predict the label of the center pixel.
-In this section, we introduce our proposed Diff-HSI in detail, including the unsupervised diffusion-based pretraining step as well as the supervised finetuning step with multi-timestep representation generation.
-III-A Diffusion-based Unsupervised Spectral-Spatial Feature Learning
-In order to learn complex spectral-spatial relations and label-agnostic information of HSI data, we pretrain a diffusion model in an unsupervised manner in the first step of our Diff-HSI.
-We first give a brief review of DDPM and introduce the basic principles to make our method easier to understand.
-Then, we explain our unsupervised feature learning procedure in detail through diffusion-based pretraining with HSI data.
+In the following section, we introduce the proposed MTMSD in detail.
+III-A Diffusion Feature Learning on unlabeled HSI
 III-A1 A Brief Review of DDPM
-Denoising diffusion probabilistic models are a class of likelihood-based models that reconstruct the distribution of training data via an encoder-decoder denoising model.
+DDPMs are a class of likelihood-based models that reconstruct the distribution of training data via an encoder-decoder denoising model.
 The denoising model is trained to remove noise from the training data destructed by Gaussian noises step-by-step.
@@
 In detail, learned variances $\sigma_{\theta}(x_{t},t)$ and an improved cosine noise schedule proposed in [25] lead to enhanced distribution learning ability.
-Figure 3: 2D t-SNE visualization of features from different timesteps on the Indian Pines dataset.
-(a) original spectral feature.
-(b) feature at $t=100$.
-(c) feature at $t=300$.
-(d) feature at $t=500$.
-(e) feature at $t=700$.
-(f) feature at $t=900$.
 III-A2 Unsupervised Hyperspectral Diffusion Pretraining
-Using the training skills and optimization objectives mentioned above, our DDPM is trained by unlabeled hyperspectral data.
-Suppose that $S$ is the set of all samples in a dataset, $D$ is a denoising diffusion probabilistic model, and $E$ is an ensemble of lightweight classifiers.
-The target of the unsupervised pretraining step is to learn label-agnostic spectral-spatial features on unlabelled data sampled from $S$ so that the features can be leveraged in the finetuning step to train ensemble lightweight classifiers using a small amount of labeled data sampled from $S_{labeled}\in S$ for classification.
-In order to learn effective spectral-spatial features in an unsupervised manner, we first pretrain a denoising diffusion probabilistic model with unlabeled patches randomly cropped from the HSI dataset.
-Before training, the data is pre-processed by principal components analysis (PCA) and patch cropping operation.
-Then, patches are randomly sampled from compressed HSI for DDPM pretraining.
-In detail, given an unlabeled patch $x_{0}\in\mathcal{R}^{H\times H}$, we gradually add Gaussian noise to the HSI patch according to the variance schedule $\{\beta_{t}\}_{t=0}^{T}$ in the diffusion process where $T$ is the total number of the timestep.
+Using the training skills and optimization objectives mentioned above, the DDPM for diffusion feature learning is trained with unlabeled hyperspectral data.
+Before training, the HSI data is pre-processed by principal components analysis (PCA) and random patch cropping operation.
+Then, given an unlabeled patch $x_{0}\in\mathcal{R}^{H\times H\times D}$, where $H$ is the patch size, and $D$ is the number of PCA components, we gradually add Gaussian noise to the unlabeled HSI patch according to the cosine variance schedule $\{\beta_{t}\}_{t=0}^{T}$ in the diffusion process, where $T$ is the total number of the timestep.
 Then, in the reverse process, a denoising U-Net is trained to predict the noise added on $x_{t-1}$ taking noisy patch $x_{t}$ and timestep $t$ as inputs.
-And $x_{t-1}$ is calculated by subtracting the predicted noise from $x_{t}$.
-In other words, the model tries to capture some useful information of $x_{t}$ to reconstruct $x_{t-1}$ that is closer to $x_{0}$.
+And the sample $x_{0}$ can be obtained from the noise patch $x_{t}$ by the iterative denoising steps according to Eq.
+3.
 In each step of the training process, the timestep $t$ is randomly sampled from 0 to $T$.
-In the pretraining procedure with unlabeled hyperspectral data, the DDPM is able to automatically learn the label-agnostic prior of HSI data that reveals the complex dependencies and relations in the spectral and spatial dimensions.
-At different timesteps, different input noisy patches and their optimization goals allow the DDPM to attend to varied information of the data, which is demonstrated to be critical to HSI classification.
-III-B Multi-Timestep Representation for Supervised Finetuning
-After unsupervised pretraining, in order to explore the rich information learned by the pretrained DDPM, the second step of our method is to extract effective features from the pretrained DDPM and construct multi-timestep representations for HSI classification.
-We first extract abundant features and construct a timestep-wise feature bank.
-Then, a dynamic feature fusion module is designed to automatically fuse the features into multi-timestep representations.
-Finally, ensemble lightweight classifiers are trained for classification.
-III-B1 Timestep-Wise Feature Bank
-In the finetuning step, the parameters of the pretrained DDPM are frozen.
-The denoising U-Net is able to capture meaningful information from the input data in the reverse process.
-Leveraging the rich and diverse information captured in the reverse steps is crucial for accurate classification.
-Therefore, timestep-wise spectral-spatial features are extracted from the intermediate hierarchies of DDPM at different timesteps to construct effective representations that contain diverse information of input HSI.
-Given an input patch $x_{0}\in\mathcal{R}^{H\times H}$ randomly cropped from HSI and a timestep $t$, Gaussian noise is gradually added to $x_{0}$ according to the variance schedule through diffusion process $q(x_{t}|x_{0})$.
-The noisy patch $x_{t}$ can be directly obtained by equation 2.
-Then, the noisy patch $x_{t}$ is fed into the pretrained denoising U-Net to obtain hierarchal features from the U-Net decoder, which implies rich spectral information as well as multi-scale spatial information, as shown in Fig.
+The U-Net denoising model $\epsilon_{\theta}(x_{t},t)$ is optimized by minimizing Eq.
+5.
+The parameters in the pretraining process, such as patch size, number of PCA components, and total pretraining steps, will be discussed in Sec IV.
+F.
+Figure 3: Purification index generation of our proposed class & timestep-oriented feature purification module.
+III-B Multi-Timestep Multi-Stage Diffusion Feature Extraction
+After diffusion feature learning on unlabeled HSI, there exist abundant multi-timestep multi-stage diffusion features that contain both contextual semantics and textual information hidden in the denoising U-Net with different timestep $t$.
+To model the complex spectral-spatial relations in HSIs, we extract multi-timestep multi-stage features from all stages of the fixed pretrained denoising U-Net decoder with all timesteps, and construct the timestep-wise center and global feature bank by center extraction and average pooling.
+Specifically, given a labeled patch $x_{0}^{{}^{\prime}}\in\mathcal{R}^{H\times H\times D}$ preprocessed by PCA to $D$ channels, $x_{0}^{{}^{\prime}}$ is corrupted by adding Gaussian noise according to Eq.
+2 and obtain $\{x^{{}^{\prime}}_{t_{i}}\}_{i=1}^{m}$ at a set of timesteps $\{t_{i}\}_{i=1}^{m}$ that are sampled from $[0,T]$ at equal intervals.
+Then, the noisy patches $\{x^{{}^{\prime}}_{t_{i}}\}_{i=1}^{m}$ are fed into the pretrained denoising U-Net to extract multi-timestep multi-stage diffusion features from all stages of the U-Net decoder.
+The different layer features are jointly upsampled to $H\times H$ and then concatenated to get the multi-stage feature $f_{t_{i}}\in\mathcal{R}^{H\times H\times d}$ at timestep $t_{i}$.
+For each multi-stage feature $f_{t_{i}}$, we only reserve the center feature $c_{t_{i}}\in\mathcal{R}^{1\times 1\times d}$ located as $(\frac{H}{2},\frac{H}{2})$ corresponding to the center pixel, and obtain the global feature $g_{t_{i}}\in\mathcal{R}^{1\times 1\times d}$ through global average pooling, which largely reduce computational cost with fewer memories.
+Following the above process, we construct the timestep-wise center feature bank $\mathcal{B}_{c}$ and the timestep-wise global feature bank $\mathcal{B}_{g}$:
+$\mathcal{B}_{c}=\{c_{t_{i}}|i\in\{1,...,m\},c_{t_{i}}=center(f_{t_{i}})\}$ (6)
+$\mathcal{B}_{g}=\{g_{t_{i}}|i\in\{1,...,m\},g_{t_{i}}=avgpool(f_{t_{i}})\}$ (7)
+III-C Class & Timestep-oriented Multi-Stage Feature Purification
+Multi-stage features, extracted from the denoising U-Net, embody abundant reconstruction information from the pretraining process of the diffusion model.
+Despite this richness, the information is not entirely aligned with HSI classification requirements, containing redundant features irrelevant to the task.
+Furthermore, these features exhibit a degree of repetition among themselves.
+Therefore, the class & timestep-oriented multi-stage purification is proposed to explore multi-stage features by selecting significant channels, aiming to remove the redundant information and reduce the computational cost.
+Before the multi-stage feature purification, we first generate the purification index using the prior of dataset-wise multi-timestep multi-stage feature bank, depicted in Fig.
+3.
+Specifically, for a $C$-category HSI classification dataset, $S^{j}$ is the training samples of category $j$, $j\in\{1,...,C\}$.
+The feature banks of samples in $S^{j}$ are averaged to get the category representative features of category $j$.
+Representing all training samples by the category representative features can significantly reduce computational overhead.
+Gathering all the representative features of each category, a representative feature matrix $M\in\mathcal{R}^{m\times C\times d}$ can be obtained,
+$M_{i,j}=\frac{1}{|S^{j}|}\sum_{x\in S^{j}}c_{t_{i}}(x)$ (8)
+where $c_{t_{i}}(x)$ is the feature at $t_{i}$ in the center feature bank of sample $x$, $|S^{j}|$ is the size of $S^{j}$.
+In order to purify the discriminative and effective channels for classification, a class-oriented significance score is proposed to evaluate the significance of each channel from inter-class relations.
+It aims to filter out channels that are highly homogenized and thus have little impact on classification by minimizing the inter-class similarity and maximizing the inter-class variance.
+For $k^{th}$ index, the class-oriented significance score $\tau_{class}^{k}$ is formed as:
+$\tau_{class}^{k}=-\alpha U_{class}^{k}+(1-\alpha)V_{class}^{k}$ (9)
+where $\alpha\in(0,1)$, $U_{class}^{k}$ denotes the inter-class similarity at index $k$ obtained by summing the average cosine similarities across classes at all the timesteps, and $V_{class}^{k}$ denotes the inter-class variance at index $k$ which is the sum of the variances across classes at all the timesteps, formed as follows.
+$U_{class}^{k}=\frac{1}{m}\frac{1}{c^{2}}\sum_{i=1}^{m}\sum_{p=1}^{c}\sum_{\begin{subarray}{c}q=1\\ q\neq p\end{subarray}}^{c}m_{i,p,k}\cdot m_{i,q,k}$ (10)
+$V_{class}^{k}=\frac{1}{m}\frac{1}{c}\sum_{i=1}^{m}\sum_{p=1}^{c}(M_{i,p,k}-\frac{1}{c}\sum_{q=1}^{c}M_{i,q,k})^{2}$ (11)
+Similarly, to preserve the diversity of features while reducing repetitive information at the timestep dimension different from the class dimension, the timestep-oriented significance score $\tau_{t}^{k}$ at index $k$ is designed to be calculated as:
+$\tau_{t}^{k}=-\beta U_{t}^{k}+(1-\beta)V_{t}^{k}$ (12)
+where $\beta\in(0,1)$, $U_{t}^{k}$ denotes the inter-timestep similarity at index $k$ obtained by summing the average cosine similarities across timesteps at all the classes, and $V_{t}^{k}$ is the inter-timestep variance at index $k$ which is the sum of the variances across classes at all the timesteps, defined as follows.
+$U_{t}^{k}=\frac{1}{c}\frac{1}{m^{2}}\sum_{i=1}^{c}\sum_{p=1}^{m}\sum_{\begin{subarray}{c}q=1\\ q\neq p\end{subarray}}^{m}M_{p,i,k}\cdot M_{q,i,k}$ (13)
+$V_{t}^{k}=\frac{1}{C}\frac{1}{m}\sum_{i=1}^{c}\sum_{p=1}^{m}(m_{p,i,k}-\frac{1}{m}\sum_{q=1}^{m}m_{q,i,k})^{2}$ (14)
+Finally, the class & timestep-oriented significance score of each index $k\in\{1,...,d\}$ is obtained as the final measurement.
+$\tau^{k}=\tau_{class}^{k}+\tau_{t}^{k}$ (15)
+We rank the indexes by the class & timestep-oriented significance score $\tau^{k}$ and generate the purification index by reserving the indexes of the top $K$ highest $\tau^{k}$, which indicates the most inter-class and inter-timestep divergence and discrimination.
+Then, multi-stage features in the timestep-wise center feature bank $\mathcal{B}_{c}$ multiply the purification index to obtain the purified timestep-wise center feature bank $\mathcal{B}_{c}^{{}^{\prime}}$.
+Figure 4: The structure of global feature-guided selective timestep network in our proposed selective timestep feature fusion.
+III-D Selective Timestep Feature Fusion
+For each labeled HSI patch $x_{0}^{{}^{\prime}}$, multi-timestep features contain textual information and contextual semantics from shallow to deep timesteps.
+Meanwhile, HSI data have different spectral characteristics for different HSI datasets, resulting in different spectral representation laws of timestep-wise features.
+Compared to the manual selection of single timestep feature in previous works, we explore the multi-timestep features and perform selective timestep feature fusion to adaptively select different timestep features integrating textual information and contextual semantics and learn comprehensive multi-timestep multi-stage selective representations.
+The whole process is illustrated in Fig.
 2.
-The features from different layers of the decoder are jointly upsampled to $H\times H$ and then concatenated to get the feature $f_{t}\in\mathcal{R}^{H\times H\times d}$ at timestep $t$.
-The above process is defined as a function $\mathcal{F}$ and formulated as follows:
-$f_{t}=\mathcal{F}(D,x_{t},t)$ (6)
-where $D$ is the pretrained denoising U-Net.
-For each feature $f_{t_{i}}\in\mathcal{R}^{H\times H\times d}$, we reserve only the vector corresponding to the center pixel indexed as $(\frac{H}{2},\frac{H}{2})$ to obtain the pixel-wise feature $h_{t_{i}}\in\mathcal{R}^{1\times 1\times d}$, which largely reduce computational cost with fewer parameters.
-Following the above process, we draw $m$ sets of features to construct the timestep-wise feature bank:
-=$\mathcal{B}=\{h_{t_{i}}|i\in\{1,...,m\},h_{t_{i}}=center(f_{t_{i}})\}$ (7)
-where $t_{1},...,t_{m}$ are sampled from $[0,T]$ at equal intervals.
-Features extracted at different timestep $t$ tend to capture different information of HSI data since the reverse process is a step-by-step denoising process that tries to predict further information conditioned on the given destructed data at every different timestep.
-To further demonstrate this, we visualize the features in 2-dimensional space using t-distributed stochastic neighbor embedding (TSNE) to show the diversity of features at different timesteps $t$.
-The visualization of the original spectral feature space is shown in Fig.
-3(a), and the features in the timestep-wise feature bank at timestep 100, 300, 500, 700, and 900 are visualized in Fig.
-3(b)-(f) respectively.
-As observed, the features in the timestep-wise feature bank show great diversity.
-Furthermore, with the increase of timestep $t$, the clusters of the features tend to be more loosely structured, and features of the same class tend to be formed into a few large clusters rather than dispersed into multiple smaller clusters.
-Features at smaller $t$ tend to capture stochastic variation, and features at larger $t$ tend to focus on high-level information, both of which are key to HSI classification with different categories of land covers.
-Thus the abundant features in the feature bank contain diverse and multi-level information of the input HSI data, exhibiting diverse properties and concerns.
-III-B2 Dynamic Feature Fusion Module
-For each input HSI patch $x_{0}$, a timestep-wise feature bank is obtained with abundant features containing varied characteristics of the patch.
-However, the features contain a wealth of information, but not all of it is essential.
-Using all the features for classification will bring high computational costs and redundant information, and manually selecting timestep-wise features from the feature bank is difficult to characterize spectral-spatial features effectively and robustly.
-Meanwhile, HSI data have different spectral characteristics for different HSI datasets, resulting in different spectral representation laws of timestep-wise features.
-Also, data on the same dataset is very different due to its position.
-Therefore, the dynamic feature fusion module is crucial to be proposed, and is capable of adaptively fusing features of different timesteps from the DDPM to learn effective multi-timestep representations with key information.
-To obtain dynamic multi-timestep representations, we design the dynamic feature fusion module to predict the soft dynamic weight of multi-timestep representations.
-The architecture of the dynamic feature fusion module is illustrated in Fig.
-2.
-Specifically, Let $L=[h_{t_{1}}...h_{t_{m}}]$ denotes the concatenated features from the timestep-wise feature bank as the input of the dynamic weight network.
-The dynamic weight network consists of two linear layers and outputs the $n$ sets of dynamic weights $w$ with a SoftMax function:
-$w$ $=\mathcal{M}(L,W)=SoftMax(a(L,W))$ (8)
-$=SoftMax(W_{2}\delta(W_{1}L))$
-where $M$ is the dynamic weight network, $W$ is the linear weight of $M$, $\delta$ is ReLU function, $w\in\mathcal{R}^{m\times 1\times 1\times n}$ and $L\in\mathcal{R}^{m\times 1\times 1\times n}$.
-Inspired by the design of the multi-head attention mechanism in [42], we generate multiple sets of dynamic weights, which allows the representations to focus on different information from different positions and spectral bands.
-Each dynamic weight determines the proportion of each feature in the corresponding representation.
-Then, the dynamic weights are used to weight the pixel-wise features by matrix multiplication to obtain $n$ multi-timestep representations:
-$r_{i}=w^{T}_{i}L$ (9)
-where $r_{i}$ is the i-th representation, $i\in{1,...,n}$.
-The multi-step representations will be concatenated into the final representation to be classified.
-Thus, the most appropriate and relevant representation is automatically learned from the feature bank without human intervention, saving time cost and reducing the impact of redundant information.
-$n$ will be ablated to find the most suitable value balancing the performance and computational cost.
-III-B3 Lightweight Ensemble Classifiers
-After getting the dynamic representation, a lightweight network is needed to predict the classification label.
-Inspired by [43], we train an ensemble of lightweight linear classifiers that takes the dynamic pixel representations as inputs and predicts the classification label of each pixel.
-Specifically, each classifier is trained independently, consisting of two hidden layers with ReLU activation and batch normalization.
-When testing a sample, the final predicted label is obtained by majority voting of the ensemble of pixel classifiers, as illustrated in Fig.
-2.
-This method brings more stability of prediction with a very small cost since the parameters of each classifier are very limited.
-IV Experiments and results
-In this section, we first describe the three used well-known HSI datasets, including the Indian Pines dataset, Pavia University dataset and Houston2018 dataset.
-The experimental setting is then introduced including evaluation metrics, a brief introduction of compared state-of-art methods, and implementation details.
-Then, we conduct quantitative experiments and ablation analysis to evaluate our proposed method.
-IV-A Datasets Description
-IV-A1 Indian Pines
-The Indian Pines dataset was acquired in 1992 over an area of Indian pines in North-Western Indiana by Airborne Visible Infrared Imaging Spectrometer (AVIRIS) sensor.
-It consists of $145\times 145$ pixels with a spatial resolution of 20 m and 220 spectral bands in the wavelength range of 400 to 2500 nm.
-There are 200 bands retained for classification (1-103, 109-149, 164-219) after removing the bands affected by noise.
-The dataset contains 10249 labeled pixels with 16 categories.
-We use 10% of the labeled samples for training and the rest for testing.
-The class name and the number of training and testing samples are listed in Table I.
-IV-A2 Pavia University
-The University of Pavia dataset was collected in 2003 by the reflective optics system imaging spectrometer (ROSIS-3) sensor over a part of the city of Pavia, Italy.
-The dataset consists of $610\times 340$ pixels with a spatial resolution of 1.3 m and 115 spectral bands in the wavelength range of 430 to 860 nm.
-103 out of 115 bands are used for classification after removing 12 noisy bands.
-The image contains a large number of background pixels and only 42776 labeled pixels divided into 9 classes, including asphalt, meadows, gravel, and so on.
-We use 5% of the labeled samples for training and the rest for testing.
-The class name and the number of training and testing samples are listed in Table II.
-IV-A3 Houston2018
-The Houston2018 dataset, identified as 2018 IEEE GRSS DFC dataset, was gathered in 2018 by the National Center for Airborne Laser Mapping (NCALM) over the University of Houston campus and its neighboring urban area, including HSI, multispectral LiDAR, and very high resolution RGB images.
-The HSI dataset consists of $601\times 2384$ pixels with a spatial resolution of 1 m and 48 spectral bands in the wavelength range of 380 to 1050 nm.
-It contains 504856 labeled pixels and 20 classes of interest.
-We use 5% of the labeled samples for training and the rest for testing.
-The class name and the number of training and testing samples are listed in Table III.
+Specifically, for the purified timestep-wise center feature bank $\mathcal{B}_{c}^{{}^{\prime}}=\{c_{t_{i}}^{{}^{\prime}}|i\in\{1,...,m\}\}$, our goal is to adaptively select the timestep of features in it for modeling spectral-spatial relations better.
+We first fuse all purified center features in $\mathcal{B}_{c}^{{}^{\prime}}$ to obtain the all-timestep feature $c_{t}\in\mathcal{R}^{1\times 1\times K}$ via an element-wise summation:
+$c_{t}=\sum_{i=1}^{m}c_{t_{i}}^{{}^{\prime}}$ (16)
+Then, the all-timestep feature $c_{t}$ is fed into a simple multi-layer perception (MLP) to obtain a compact feature $z\in\mathcal{R}^{1\times 1\times K_{r}}$ with fewer channels for better efficiency:
+$z=\mathcal{F}_{mlp}(c_{t})=W_{1}(\delta(BN(W_{2}c_{t})))$ (17)
+where $\delta$ is the ReLU function, $BN$ is the batch normalization, $W_{1}\in\mathcal{R}^{K\times K_{r}}$, $W_{2}\in\mathcal{R}^{K_{r}\times K_{r}}$, $K_{r}=K/2$.
+However, adaptive timestep selection solely based on timestep-wise center features leads to imprecise choices due to the absence of spatial information from neighboring pixels around the central pixel.
+Hence, we propose a global feature-guided selective timestep network, designed to incorporate global features for guiding multi-timestep feature selection, thereby enriching spatial information in modeling spatial distributions.
+And the network is depicted in Fig.
+4.
+Specifically, given a timestep-wise global feature bank $\mathcal{B}_{g}=\{g_{t_{i}}|i\in\{1,...,m\}\}$, for each purified center feature $c_{t_{i}}^{{}^{\prime}}$, the corresponding global feature $g_{t_{i}}$ is concatenated with $z$ with the transformation $\mathcal{F}_{i}$ to obtain $h_{i}$.
+Then the linear projection $\mathcal{W}_{i}$ is applied to the $h_{i}$ respectively, $i\in\{1,...,m\}$.
+To adaptively select different timesteps of center features, a softmax operator on the channel-wise digits is used to obtain the selective weights $\{w_{i}\}_{i=1}^{m}$ guided by the compact feature $z$ and the corresponding global information:
+$w^{{}^{\prime}}_{i}=\mathcal{W}(h_{i})=\mathcal{W}([\mathcal{F}_{i}(z),g_{t_{i}}])$ (18)
+$w^{c}_{i}=\frac{e^{w^{{}^{\prime}c}_{i}}}{\sum_{j=1}^{m}e^{w^{{}^{\prime}c}_{j}}}$ (19)
+where $\{\mathcal{F}_{i}\}_{i=1}^{m}$ and $\{\mathcal{W}_{i}\}_{i=1}^{m}$ are linear projections to align the channel dimension to $K$ channels, and $w^{c}_{i}$ is the $c$-th element of the selective weight $w_{i}$, $c\in\{1,...,K\}$.
+Finally, the multi-timestep multi-stage selective representation $r_{s}$ is selected and fused through the selective weights $\{w_{i}\}_{i=1}^{m}$ on each purified center feature $c_{t_{i}}^{{}^{\prime}}$:
+$r_{s}=\sum_{i=1}^{m}w_{i}c_{t_{i}}^{{}^{\prime}}$ (20)
 TABLE I: Land-cover types, the number of labeled training samples and testing samples of the Indian Pines dataset.
@@
 Total 1024 9225
-TABLE II: Land-cover types, the number of labeled training samples and testing samples of the Pavia University dataset.
+TABLE II: Land-cover types, the number of labeled training samples and testing samples of the PaviaU dataset.
 Class Land Cover Type Training Testing
@@
 Total 2138 40638
+After obtaining the multi-timestep multi-stage selective representation, a lightweight network is needed to predict the classification label.
+Inspired by [46], we train an ensemble of lightweight linear classifiers that takes the dynamic pixel representations as inputs and predicts the classification label of each pixel.
+Specifically, each classifier is trained independently, consisting of two hidden layers with ReLU activation and batch normalization.
+When testing a sample, the final predicted label is obtained by majority voting of the ensemble of pixel classifiers, as illustrated in Fig.
+2.
+This method brings more stability of prediction with a very small cost since the parameters of each classifier are very limited.
+IV Experiments and results
+In this section, we first describe four well-known HSI datasets, including the Indian Pines dataset, the Pavia University dataset, the Houston 2018 dataset, and the WHU-Hi-Longkou dataset.
+The experimental setting is then introduced including evaluation metrics, a brief introduction of compared state-of-art methods, and implementation details.
+Then, we conduct quantitative experiments and ablation analysis to evaluate our proposed method.
+IV-A Datasets Description
+IV-A1 Indian Pines
+The Indian Pines dataset was acquired in 1992 over an area of Indian pines in North-Western Indiana by Airborne Visible Infrared Imaging Spectrometer (AVIRIS) sensor.
+It consists of $145\times 145$ pixels with a spatial resolution of 20 m and 220 spectral bands in the wavelength range of 400 to 2500 nm.
+There are 200 bands retained for classification (1-103, 109-149, 164-219) after removing the bands affected by noise.
+The dataset contains 10249 labeled pixels with 16 categories.
+We use 10% of the labeled samples for training and the rest for testing.
+The class name and the number of training and testing samples are listed in Table I.
+IV-A2 Pavia University
+The Pavia University (PaviaU) dataset was collected in 2003 by the reflective optics system imaging spectrometer (ROSIS-3) sensor over a part of the city of Pavia, Italy.
+The dataset consists of $610\times 340$ pixels with a spatial resolution of 1.3 m and 115 spectral bands in the wavelength range of 430 to 860 nm.
+103 out of 115 bands are used for classification after removing 12 noisy bands.
+The image contains a large number of background pixels, and only 42776 labeled pixels are divided into 9 classes, including asphalt, meadows, gravel, and so on.
+We use 5% of the labeled samples for training and the rest for testing.
+The class name and the number of training and testing samples are listed in Table II.
+IV-A3 Houston 2018
+The Houston 2018 dataset, identified as the 2018 IEEE GRSS DFC dataset, was gathered in 2018 by the National Center for Airborne Laser Mapping (NCALM) over the University of Houston campus and its neighboring urban area, including HSI, multispectral LiDAR, and very high-resolution RGB images.
+The HSI dataset consists of $601\times 2384$ pixels with a spatial resolution of 1 m and 48 spectral bands in the wavelength range of 380 to 1050 nm.
+It contains 504856 labeled pixels and 20 classes of interest.
+We use 5% of the labeled samples for training and the rest for testing.
+The class name and the number of training and testing samples are listed in Table III.
+IV-A4 WHU-Hi-Longkou
+The WHU-Hi-Longkou dataset was acquired in 2018 by an 8-mm focal length Headwall Nano-Hyperspec imaging sensor equipped on a DJ-innovations Matrice 600 Pro UAV platform.
+It consists of $550\times 400$ pixels with a spatial resolution of 0.463 m and 270 spectral bands in the wavelength range of 400 to 1000 nm.
+It contains 204542 labeled samples and 9 object classes.
+We use 0.5% of the labeled samples for training and the rest for testing.
+The class name and the number of training and testing samples are listed in Table IV.
 TABLE III: Land-cover types, the number of labeled training samples and testing samples of the Houston 2018 dataset.
@@
 Total 25242 479614
-TABLE IV: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Indian Pines dataset.
-The best one is shown in bold.
-Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
-2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
-1 65.85 58.54 94.14 63.00 95.12 97.56 72.97 99.10 99.51
-2 99.77 76.19 97.84 92.35 97.67 98.05 88.50 96.10 98.89
-3 81.66 77.64 97.54 86.86 98.87 98.66 87.20 95.39 99.84
-4 96.71 52.11 90.70 88.96 91.55 95.31 84.90 97.24 99.81
-5 85.75 93.56 97.75 92.49 96.32 95.17 90.28 94.12 98.85
-6 97.87 98.17 99.24 99.12 99.54 99.85 97.97 99.25 100.00
-7 100.00 36.00 81.60 52.50 100.00 100.00 56.52 88.46 98.40
-8 100.00 98.60 100.00 99.16 100.00 100.00 99.48 100.00 100.00
-9 50.00 55.56 74.44 41.18 88.89 100.00 87.50 94.44 100.00
-10 35.54 82.86 94.77 93.16 97.71 94.29 86.80 95.84 98.88
-11 88.01 90.45 98.87 92.27 98.69 99.37 96.68 99.29 99.63
-12 98.13 62.55 97.83 85.44 98.13 96.63 80.83 93.37 98.91
-13 99.46 88.65 99.24 99.02 97.28 100.00 100.00 100.00 99.89
-14 99.91 99.39 99.18 96.73 99.91 97.89 99.90 95.25 100.00
-15 91.35 86.17 93.95 83.41 98.84 97.12 96.80 99.15 99.25
-16 86.90 45.24 98.33 93.50 95.54 94.05 84.00 100.00 98.10
-OA (%) 87.77 85.42 97.75 92.31 97.47 97.95 92.69 97.02 99.46
-AA (%) 86.06 75.10 94.71 84.95 96.57 97.75 88.15 96.00 99.37
-$\kappa$ 0.8603 0.8324 0.9743 0.9124 0.9711 0.9766 0.9162 0.9660 0.9938
-TABLE V: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$, as well as the accuracies for each class on the Pavia University dataset.
-The best one is shown in bold.
-Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
-2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
-1 99.68 97.02 98.81 96.21 99.33 99.38 94.20 99.62 100.00
-2 99.41 99.97 99.83 99.64 99.92 99.80 99.58 99.98 99.99
-3 86.56 92.98 92.45 87.65 98.29 98.35 78.83 91.02 100.00
-4 98.18 97.53 98.32 96.64 98.49 99.52 97.53 98.40 99.51
-5 99.84 99.06 99.65 99.97 99.53 100.00 100.00 100.00 100.00
-6 100.00 99.10 99.43 99.56 100.00 99.75 95.30 98.32 100.00
-7 99.84 79.10 99.76 90.30 99.13 99.60 97.42 99.61 100.00
-8 100.00 97.34 99.32 94.60 98.05 98.63 96.87 98.36 99.82
-9 98.22 95.22 99.82 98.49 95.44 99.33 98.61 99.89 99.96
-OA (%) 98.86 97.88 99.10 97.54 99.21 99.53 96.77 99.02 99.94
-AA (%) 97.97 95.26 98.60 95.88 98.69 99.37 95.37 98.36 99.92
-$\kappa$ 0.9848 0.9719 0.9881 0.9674 0.9915 0.9937 0.9571 0.9870 0.9993
-TABLE VI: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Houston2018 dataset.
-The best one is shown in bold.
-Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
-2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
-1 84.33 82.02 86.30 92.36 79.93 79.50 91.60 88.99 89.41
-2 95.98 96.64 95.32 95.08 93.44 96.55 93.92 97.53 96.24
-3 96.92 96.00 99.72 96.21 99.66 100.00 97.60 99.23 99.88
-4 98.13 95.55 97.49 98.48 96.64 97.62 95.92 97.98 99.26
-5 91.26 79.16 86.09 91.87 90.11 95.01 85.79 92.51 96.86
-6 99.91 97.51 98.48 99.62 99.62 99.91 98.50 99.58 100.00
-7 80.63 71.94 93.68 27.01 85.45 95.65 61.15 96.85 97.63
-8 94.29 88.62 91.62 95.42 98.72 99.18 91.15 94.20 99.62
-9 99.54 92.80 97.88 98.60 99.09 99.35 95.12 99.07 99.63
-10 95.44 71.61 81.72 88.22 91.15 92.64 76.72 88.61 95.89
-11 77.86 73.17 69.44 27.01 80.97 85.73 70.22 82.28 93.02
-12 15.26 11.10 0.51 31.46 41.69 33.43 4.79 42.76 50.23
-13 87.30 69.12 84.59 91.79 94.49 96.79 90.52 93.26 97.72
-14 94.97 96.18 89.65 92.91 96.97 99.17 87.23 97.02 99.23
-15 95.40 98.98 99.34 99.33 99.30 99.92 99.06 99.59 99.81
-16 93.62 90.40 91.00 96.36 97.84 98.15 92.01 96.36 99.69
-17 86.33 20.86 0.00 22.78 69.21 90.65 0.00 100.00 94.10
-18 96.64 89.05 93.66 91.61 93.07 97.85 90.43 94.24 99.46
-19 99.24 95.45 96.92 96.53 97.97 99.84 96.09 99.59 99.89
-20 92.77 93.92 99.12 99.77 99.96 100.00 97.27 99.89 99.94
-OA (%) 94.84 86.88 91.61 90.65 95.48 96.69 90.34 95.38 98.07
-AA (%) 88.79 80.50 82.63 80.75 90.26 92.85 80.76 92.98 95.37
-$\kappa$ 0.9325 0.8313 0.8906 0.8784 0.9412 0.9570 0.8751 0.9398 0.9748
-IV-B Experimental Setting
-IV-B1 Evaluation Metrics
-We evaluate the performance of all methods by three widely used indexes: overall accuracy (OA), average accuracy (AA), and Kappa coefficient ($\kappa$).
-IV-B2 Comparison with State-of-the-art Methods
-To demonstrate the effectiveness of our proposed method, we compare our classification performance with several state-of-the-art approaches, including CNN-based and transformer-based supervised methods, and unsupervised methods.
-We use the most effective setting for each of these methods.
-  • The 2-D CNN architecture contains three 2-D convolution blocks and a softmax layer.
-Each convolution block consists of a 2-D convolution layer, a BN layer, an avg-pooling layer, and a ReLU activation function.
-  • The 3-D CNN contains three 3-D convolution blocks and a softmax layer.
-Each 3-D convolution block consists of a 3-D convolution layer, a BN layer, a ReLU activation function, and a 3-D convolution layer with step size 2.
-  • The SSRN[14] is a spectral-spatial residual network based on 3-D CNN and residual connection.
-Spatial residual blocks and spatial residual blocks are designed to extract discriminative features from HSI data.
-  • For SF[16], there are four encoder blocks in the ViT-based network.
-In the transformer framework of SF, group-wise spectral embedding and cross-layer adaptive fusion modules are adopted to capture local spectral representations from neighboring bands.
-  • The SSFTT[17] systematically combines CNN network and transformer structure to exploit spectral-spatial information in the HSI, with a Gaussian weighted feature tokenizer module making the samples more separable.
-  • The GAHT[18] is a end-to-end group-aware transformer method with three-stage hierarchical framework.
-  • The 3DCAE[21] is an unsupervised method using the encoder-decoder backbone with 3D convolution operation to learn spectral-spatial features.
-  • The UMSDFL[22] is an unsupervised method using encoder and decoder with convolutional layers to learn spectral-spatial features.
-A clustering branch and a multi-layer fusion module is designed to enhance the features.
-IV-B3 Implementation Details
-The proposed Diff-HSI was implemented using the Pytorch framework.
-The patch size is set to $48\times 48$, and the dimension of PCA is set to 10.
-In the diffusion-pretraining procedure, we use Kullback-Leibler Divergence Loss as the loss function.
-And the Adam optimizer is adopted with a batch size of 128 and a learning rate of 1e-4.
-In the second stage, the cross-entropy loss is used in lightweight classifiers.
-$m$, $n$, and $k$ are set to be 20, 3, and 5, respectively.
-We adopt the Adam Optimizer and the Cosine Annealing as our training schedule.
-The original learning rate and minimum learning rate are set to be 1e-4 and 5e-6, respectively.
-The number of epochs is set to 100 for all datasets.
-Fairly, we calculate the results by averaging the results of ten repeated experiments with different training sample selections.
-IV-C Quantitative Results and Analysis
-Quantitative classification results in terms of class-specific accuracy, OA, AA, and $\kappa$ of the compared methods on the Indian Pines, PaviaU, and Houston 2018 datasets are listed in Table IV, V, and VI, respectively.
-And the classification maps of all methods are shown in Fig.
-4, 5 and 6.
-Compared with other methods, our proposed Diff-HSI achieves the highest OA, AA, and $\kappa$ on the three datasets.
-According to the results, the CNN-based supervised methods obtain good performance owing to the ability of capturing local spatial information.
-Besides, since transformers are capable of capturing sequential information, transformer-based supervised methods also achieve competitive performance.
-The GAHT method achieves the best results in CNN-based and transformer-based supervised methods because it balances the local information and the global dependencies.
-However, the methods that learn explicit mappings have bottlenecks in learning label-agnostic information and complex relations, which limits performance.
-Some unsupervised methods are proposed to tackle the problem by learning representative features through an encoder-decoder network.
-Limited by the model architecture, the features learned are not discriminative and rich enough for HSI classification lacking high-level information.
-Therefore, the performance of these methods is even lower than that of some explicit learning methods.
-Our proposed Diff-HSI introduces the diffusion model to HSI classification, which has the powerful ability to model complex relations.
-Furthermore, the features learned by our method are of great diversity, enriching the representation capability.
-Thus, our Diff-HSI outperforms all the previous methods on the three datasets.
-Notably, the classification performance of the Houston2018 Dataset is largely improved compared with the previous SOTA method in terms of OA (98.07% versus 96.69%), AA (95.37% versus 92.98%), and $\kappa$ (0.9748 versus 0.9570), which especially demonstrate our effectiveness.
-((a))
-((b))
-((c))
-((d))
-((e))
-((f))
-((g))
-((h))
-((i))
-((j))
+TABLE IV: Land-cover types, the number of labeled training samples and testing samples of the WHU-Hi-Longkou dataset.
+Class Land Cover Type Training Testing
+1 Corn 172 34339
+2 Cotton 42 8332
+3 Sesame 15 3016
+4 Broad-leaf soybean 316 62896
+5 Narrow-leaf soybean 21 4130
+6 Rice 59 11795
+7 Water 335 66721
+8 Roads and houses 36 7088
+9 Mixed weed 26 5203
+Total 1022 203520
+TABLE V: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Indian Pines dataset.
+The best results are shown in bold.
+Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
+1 65.85 58.54 94.14 63.00 95.12 97.56 72.97 100.00 99.10 100.00 100.00
+2 99.77 76.19 97.84 92.35 97.67 98.05 88.50 89.34 96.10 97.90 99.52
+3 81.66 77.64 97.54 86.86 98.87 98.66 87.20 95.98 95.39 98.93 99.01
+4 96.71 52.11 90.70 88.96 91.55 95.31 84.90 95.31 97.24 100.00 99.62
+5 85.75 93.56 97.75 92.49 96.32 95.17 90.28 88.74 94.12 94.02 98.62
+6 97.87 98.17 99.24 99.12 99.54 99.85 97.97 99.09 99.25 99.54 99.97
+7 100.00 36.00 81.60 52.50 100.00 100.00 56.52 56.00 88.46 100.00 99.20
+8 100.00 98.60 100.00 99.16 100.00 100.00 99.48 99.77 100.00 100.00 100.00
+9 50.00 55.56 74.44 41.18 88.89 100.00 87.50 100.00 94.44 100.00 100.00
+10 35.54 82.86 94.77 93.16 97.71 94.29 86.80 87.77 95.84 98.51 98.79
+11 88.01 90.45 98.87 92.27 98.69 99.37 96.68 96.88 99.29 99.77 99.67
+12 98.13 62.55 97.83 85.44 98.13 96.63 80.83 83.71 93.37 91.76 98.84
+13 99.46 88.65 99.24 99.02 97.28 100.00 100.00 100.00 100.00 99.46 99.78
+14 99.91 99.39 99.18 96.73 99.91 97.89 99.90 99.30 95.25 99.91 99.86
+15 91.35 86.17 93.95 83.41 98.84 97.12 96.80 98.56 99.15 98.27 99.42
+16 86.90 45.24 98.33 93.50 95.54 94.05 84.00 97.62 100.00 98.81 98.10
+OA (%) 87.77 85.42 97.75 92.31 97.47 97.95 92.69 94.34 97.02 98.54 99.45
+AA (%) 86.06 75.10 94.71 84.95 96.57 97.75 88.15 93.00 96.00 98.56 99.40
+$\kappa$ 0.8603 0.8324 0.9743 0.9124 0.9711 0.9766 0.9162 0.9353 0.9660 0.9833 0.9937
 Figure 5: Classification maps obtained by different methods on the Indian Pines dataset.
@@
 (h) 3DCAE (OA=92.69%).
-(i) UMSDFL (OA=97.02%).
-(j) Diff-HSI (OA=99.46%).
-((a))
-((b))
-((c))
-((d))
-((e))
-((f))
-((g))
-((h))
-((i))
-((j))
-Figure 5: Classification maps obtained by different methods on the Pavia University dataset.
+(i) 3DAES (OA=94.34%).
+(j) UMSDFL (OA=97.02%).
+(k) SpectralDiff (OA=98.54%).
+(l) MTMSD (OA=99.45%).
+TABLE VI: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$, as well as the accuracies for each class on the PaviaU dataset.
+the best results are shown in bold.
+Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
+1 99.68 97.02 98.81 96.21 99.33 99.38 94.20 97.94 99.62 99.89 100.00
+2 99.41 99.97 99.83 99.64 99.92 99.80 99.58 99.00 99.98 100.00 100.00
+3 86.56 92.98 92.45 87.65 98.29 98.35 78.83 89.57 91.02 98.75 100.00
+4 98.18 97.53 98.32 96.64 98.49 99.52 97.53 98.63 98.40 97.35 99.59
+5 99.84 99.06 99.65 99.97 99.53 100.00 100.00 100.00 100.00 95.93 100.00
+6 100.00 99.10 99.43 99.56 100.00 99.75 95.30 97.11 98.32 100.00 100.00
+7 99.84 79.10 99.76 90.30 99.13 99.60 97.42 94.77 99.61 100.00 100.00
+8 100.00 97.34 99.32 94.60 98.05 98.63 96.87 98.20 98.36 99.77 99.87
+9 98.22 95.22 99.82 98.49 95.44 99.33 98.61 97.44 99.89 94.44 99.79
+OA (%) 98.86 97.88 99.10 97.54 99.21 99.53 96.77 97.92 99.02 99.46 99.95
+AA (%) 97.97 95.26 98.60 95.88 98.69 99.37 95.37 96.96 98.36 98.46 99.92
+$\kappa$ 0.9848 0.9719 0.9881 0.9674 0.9915 0.9937 0.9571 0.9725 0.9870 0.9929 0.9994
+Figure 6: Classification maps obtained by different methods on the PaviaU dataset.
 (a) Ground truth.
@@
 (h) 3DCAE (OA=96.77%).
-(i) UMSDFL (OA=99.02%).
-(j) Diff-HSI (OA=99.94%).
-((a))
-((b))
-((c))
-((d))
-((e))
-((f))
-((g))
-((h))
-((i))
-((j))
+(i) 3DAES (OA=97.92%).
+(j) UMSDFL (OA=99.02%).
+(k) SpectralDiff (OA=99.46%).
+(l) MTMSD (OA=99.95%).
+IV-B Experimental Setting
+IV-B1 Evaluation Metrics
+We evaluate the performance of all methods by three widely used indexes: overall accuracy (OA), average accuracy (AA), and Kappa coefficient ($\kappa$).
+IV-B2 Comparison with State-of-the-art Methods
+To demonstrate the effectiveness of our proposed method, we compare our classification performance with several state-of-the-art approaches using the most effective setting for these methods.
+  • The 2-D CNN[47] architecture contains three 2-D convolution blocks and a softmax layer.
+Each convolution block consists of a 2-D convolution layer, a BN layer, an avg-pooling layer, and a ReLU activation function.
+  • The 3-D CNN[47] contains three 3-D convolution blocks and a softmax layer.
+Each 3-D convolution block consists of a 3-D convolution layer, a BN layer, a ReLU activation function, and a 3-D convolution layer with step size 2.
+  • The SSRN[14] is a spectral-spatial residual network based on 3-D CNN and residual connection.
+Spatial residual blocks and spatial residual blocks are designed to extract discriminative features from HSI data.
+  • For SF[16], group-wise spectral embedding and cross-layer adaptive fusion modules in the transformer framework are adopted to capture local spectral representations from neighboring bands.
+  • The SSFTT[17] systematically combines CNN network and transformer structure to exploit spectral-spatial information in the HSI, with a Gaussian weighted feature tokenizer module making the samples more separable.
+  • The GAHT[18] is a end-to-end group-aware transformer method with three-stage hierarchical framework.
+  • The 3DCAE[20] is an unsupervised method using an encoder-decoder backbone with 3D convolution operation to learn spectral-spatial features.
+  • The 3DAES[48] is a semi-supervised method using an autoencoder to extract spectral-spatial features from unlabeled samples and then optimizes the siamese network and classifier using constructed sample pairs.
+  • The UMSDFL[21] is an unsupervised method using encoder and decoder with convolutional layers to learn spectral-spatial features.
+A clustering branch and a multi-layer fusion module are designed to enhance the features.
+  • The SpectralDiff[32] is a diffusion-based unsupervised method that learns diffusion features through the spectral-spatial diffusion module and feeds diffusion features into the attention-based classification module.
+IV-B3 Implementation Details
+The proposed MTMSD was implemented using the Pytorch framework.
+The patch size is set to $48\times 48$, and the dimension of PCA is set to $8/N$, $N$ is the number of spectral bands of the dataset.
+In the diffusion-pretraining procedure, we use Kullback-Leibler Divergence Loss as the loss function.
+And the Adam optimizer is adopted with a batch size of 128 and a learning rate of 1e-4, training a total of 40k steps.
+In the feature-exploring stage, the cross-entropy loss is used in lightweight classifiers.
+$m$ and $K$ are set to be 19 and 5, respectively.
+We adopt the Adam Optimizer and the Cosine Annealing as our training schedule.
+The original learning rate and minimum learning rate are set to be 1e-4 and 5e-6, respectively.
+The number of epochs is set to 100 for all datasets.
+We calculate the results fairly by averaging the results of ten repeated experiments with different training sample selections.
+IV-C Quantitative Results and Analysis
+IV-C1 Classification Results Compared with SOTA Methods
+Quantitative classification results in terms of class-specific accuracy, OA, AA, and $\kappa$ of the compared methods on the Indian Pines, PaviaU, Houston 2018 and Longkou datasets are listed in Table V, VI, VII and VIII, respectively.
+And the classification maps of all methods are shown in Fig.
+5, 6, 7 and 8.
+TABLE VII: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Houston 2018 dataset.
+the best results are shown in bold.
+Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
+1 87.12 82.02 86.30 92.36 79.93 79.50 91.60 81.80 88.99 83.47 88.87
+2 92.05 96.64 95.32 95.08 93.44 96.55 93.92 93.30 97.53 92.48 96.29
+3 96.92 96.00 99.72 96.21 99.66 100.00 97.60 89.23 99.23 99.38 99.93
+4 98.05 95.55 97.49 98.48 96.64 97.62 95.92 94.39 97.98 97.01 99.32
+5 87.25 79.16 86.09 91.87 90.11 95.01 85.79 81.95 92.51 86.25 97.01
+6 95.36 97.51 98.48 99.62 99.62 99.91 98.50 97.72 99.58 99.58 100.00
+7 96.44 71.94 93.68 27.01 85.45 95.65 61.15 83.40 96.85 87.35 97.40
+8 97.15 88.62 91.62 95.42 98.72 99.18 91.15 91.27 94.20 98.73 99.81
+9 98.24 92.80 97.88 98.60 99.09 99.35 95.12 97.12 99.07 99.31 99.72
+10 93.88 71.61 81.72 88.22 91.15 92.64 76.72 76.90 88.61 87.73 96.63
+11 75.85 73.17 69.44 27.01 80.97 85.73 70.22 71.14 82.28 79.33 93.96
+12 12.55 11.10 0.51 31.46 41.69 33.43 4.79 3.19 42.76 37.24 55.30
+13 85.24 69.12 84.59 91.79 94.49 96.79 90.52 84.62 93.26 95.64 97.87
+14 77.12 96.18 89.65 92.91 96.97 99.17 87.23 95.41 97.02 98.09 99.27
+15 94.45 98.98 99.34 99.33 99.30 99.92 99.06 97.31 99.59 99.47 99.92
+16 93.43 90.40 91.00 96.36 97.84 98.15 92.01 87.79 96.36 98.42 99.84
+17 64.75 20.86 0.00 22.78 69.21 90.65 0.00 46.04 100.00 84.17 100.00
+18 91.70 89.05 93.66 91.61 93.07 97.85 90.43 83.89 94.24 93.91 99.60
+19 96.88 95.45 96.92 96.53 97.97 99.84 96.09 94.55 99.59 98.47 99.99
+20 99.83 93.92 99.12 99.77 99.96 100.00 97.27 96.30 99.89 100.00 100.00
+OA (%) 93.38 86.88 91.61 90.65 95.48 96.69 90.34 90.39 95.38 95.28 98.29
+AA (%) 86.71 80.50 82.63 80.75 90.26 92.85 80.76 82.37 92.98 90.80 96.04
+$\kappa$ 0.9137 0.8313 0.8906 0.8784 0.9412 0.9570 0.8751 0.8748 0.9398 0.9385 0.9777
 Figure 7: Classification maps obtained by different methods on the Houston 2018 dataset.
@@
 (h) 3DCAE (OA=90.34%).
-(i) UMSDFL (OA=95.38%).
-(j) Diff-HSI (OA=98.07%).
+(i) 3DAES (OA=90.39%).
+(j) UMSDFL (OA=95.38%).
+(k) SpectralDiff (OA=95.28%).
+(l) MTMSD (OA=98.29%).
+TABLE VIII: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$, as well as the accuracies for each class on the WHU-Hi-Longkou dataset.
+the best results are shown in bold.
+Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
+1 99.44 99.55 99.80 99.76 99.85 99.91 99.71 98.85 99.93 99.53 99.97
+2 95.44 93.18 98.82 89.20 95.96 98.99 95.07 98.78 97.40 97.49 99.67
+3 81.76 93.83 91.11 97.25 93.50 96.68 85.18 80.64 91.25 100.00 97.67
+4 99.99 98.70 99.71 98.42 99.04 99.55 98.26 99.71 99.13 99.46 99.81
+5 75.45 83.74 94.04 83.80 92.42 95.96 60.39 84.02 94.19 96.39 97.31
+6 98.66 98.88 99.89 97.69 99.34 99.77 99.49 98.76 99.76 99.33 99.86
+7 99.96 99.99 99.97 99.98 99.99 99.99 99.99 99.94 99.99 99.85 99.99
+8 98.41 96.66 96.33 89.45 97.19 96.46 96.67 95.40 97.95 86.82 96.77
+9 91.08 94.77 93.95 73.44 96.16 86.51 92.89 89.54 88.93 87.28 96.24
+OA (%) 98.58 98.48 99.28 97.47 99.02 99.19 97.86 98.54 98.99 98.71 99.61
+AA (%) 93.36 95.37 97.07 92.11 97.05 97.09 91.96 93.96 96.20 96.24 98.59
+$\kappa$ 0.9812 0.9801 0.9905 0.9668 0.9872 0.9893 0.9718 0.9807 0.9868 0.9830 0.9949
+Compared with other methods, our proposed MTMSD achieves the highest OA, AA, and $\kappa$ on four datasets.
+According to the results, the CNN-based supervised methods, 2-D CNN, 3-D CNN, and SSRN, obtain good performance owing to their ability to capture local spatial information.
+Besides, since transformers are capable of capturing sequential information, transformer-based supervised methods, SF, SSFTT, and GAHT, also achieve competitive performance.
+Unsupervised methods, 3DCAE, 3DAES, and UMSDFL are proposed to tackle the problem of limited samples by learning representative features without any labeled samples.
+Limited by the model architecture, the features learned in an unsupervised manner are not discriminative enough for HSI classification lacking high-level information.
+Therefore, the performance of 3DCAE and 3DAES is even lower than that of some explicit learning methods.
+SpetralDiff introduces the diffusion model to HSI classification and achieves competitive performance due to the advantages of diffusion features.
+However, the feature used in the method is from a single timestep and a single layer with the loss of some key information, which limits the performance.
+Our proposed MTMSD explores the value of multi-timestep multi-stage diffusion features through class & timestep-oriented multi-stage feature purification and selective timestep feature fusion, effectively modeling complex spectral-spatial relations due to the adaptive integration of contextual semantics and textual details.
+Thus, our MTMSD outperforms all the previous methods on four datasets: Indian Pines, PaviaU, Houston 2018, and Longkou.
+Notably, the classification performance of the Houston 2018 Dataset is largely improved compared with the previous SOTA method in terms of OA (98.29% versus 96.69%), AA (96.04% versus 92.98%), and $\kappa$ (0.9777 versus 0.9570), which especially demonstrate our effectiveness.
+Figure 8: Classification maps obtained by different methods on the WHU-Hi-Longkou dataset.
+(a) Ground truth.
+(b) 2-D CNN (OA=98.58%).
+(c) 3-D CNN (OA=98.48%).
+(d) SSRN (OA=99.28%).
+(e) SF (OA=97.47%).
+(f) SSFTT (OA=99.02%).
+(g) GAHT (OA=99.19%).
+(h) 3DCAE (OA=97.86%).
+(i) 3DAES (OA=98.54%).
+(j) UMSDFL (OA=98.99%).
+(k) SpectralDiff (OA=98.71%).
+(l) MTMSD (OA=99.61%).
+IV-C2 Classification Results with Different Proportions of Training Samples
+The classification results with different proportions of training samples are shown in Fig.
+9.
+It can be observed that the performance increases with the percentages of training samples.
+Our method outperforms the compared method consistently in terms of OA on four datasets.
+Especially, using only 2% of the training sample, our method achieves comparable results to other methods using 5% of the training sample on the Houston 2018 dataset.
 IV-D Ablation Studies
 In this section, we analyze the effect of the components in our method.
-IV-D1 Ablation for Set Number of Dynamic Weights
-Our dynamic feature fusion module contains $n$ sets of dynamic weights which impact the classification performance.
-The change of results by changing $n$ is shown in Tabel VII.
-We reveal that the best performance is achieved when $n$ is taken as 3.
-With smaller $n$, only part of the useful information is obtained from the feature bank, which is insufficient for classification.
-On the other hand, larger $n$ will bring more parameters and excess information.
-IV-D2 Ablation for Pretraining Steps
-We validate the effectiveness of pretraining on the Indian Pines dataset in terms of OA, AA, and $\kappa$.
-As shown in Table VIII, only 10k steps pretraining brings dramatic improvement (more than 30% OA) to the final classification performance.
-Furthermore, as the pretraining steps increase, the performance continues to rise to the best at around 40k steps.
-IV-D3 Sensitivity Analysis of Timestep T
+IV-D1 Ablation for Class & Timestep-oriented Multi-Stage Feature Purification
+Table IX presents the effect of class & timestep-oriented multi-stage feature purification (CTMSFP) across the four datasets.
+The multi-stage features extracted from the diffusion model are not entirely aligned with the HSI classification task.
+Therefore, the CTMSFP is proposed to reduce the redundancy of the features and maintain efficiency.
+All experiments are conducted on an RTX 3090 GPU, with a consistent batch size of 64 used during training.
+As shown in Table IX, the results demonstrate that CTMSFP significantly reduces the number of parameters and the GPU memory consumption during training.
+This reduction is due to the channel-wise purification performed by CTMSFP before the features are input into the model, which substantially decreases the size of both the input tensor and the model structure.
+Furthermore, the average inference time is reduced by 25%, effectively enhancing the model’s computational efficiency, while still ensuring a slight performance improvement.
+We also compare the performance and efficiency of another diffusion-based HSI classification model, SpectralDiff, which utilizes single-timestep single-stage diffusion features.
+Our method performs independent linear transformations on features from different timesteps, allowing tailored transformations for diverse multi-timestep feature patterns, resulting in a larger number of parameters.
+However, our method with CTMSFP significantly leads to performance, inference speed, and GPU memory consumption.
+((a))
+((b))
+((c))
+((d))
+Figure 9: Classification performance of the compared methods with different proportions of training samples on four datasets. (a) Indian Pines. (b) PaviaU. (c) Houston 2018. (d) Longkou.
+TABLE IX: Ablation for Class & Timestep-oriented Multi-Stage Feature Purification (CTMSFP) on four datasets in terms of OA, inference time (IT), parameters and GPU memory.
+The best result are shown in bold.
+Method Indian Pines PaviaU Houston 2018 Longkou Param. (M) GPU memory (G)
+OA (%) IT (s) OA (%) IT (s) OA (%) IT (s) OA (%) IT (s)
+SpectralDiff 98.54 10.92 99.46 22.41 95.28 268.88 98.71 178.25 1.38 3.09
+MTMSD (w/o CTMSFP)
+99.39 4.55 99.90 22.78 98.20 215.20 99.51 85.14 55.91 1.28
+MTMSD 99.45 3.52 99.95 12.41 98.29 168.52 99.61 61.53 20.18 0.46
+IV-D2 Ablation for Selective Timestep Feature Fusion
+The ablation results for the selective timestep feature fusion are shown in Table.
+X.
+In the manual selection method, we extract features from a single timestep for classification.
+The optimal timestep is selected for each dataset, and the best result is shown in the table.
+However, using the features from only one single timestep leads to the loss of abundant spectral-spatial information.
+Although the optimal timestep is chosen for each dataset, it lacks adequate information, only containing textural features or semantics to model spectral-spatial relations, and is not flexible enough to accommodate different patch data, both of which limit performance.
+The average fusion method considers features from different timestep $t$, obtaining better results than manual selection.
+However, assigning a uniform selection weight to features across all the timesteps indiscriminately does not fully harness their potential because the significance of features from different $t$ for different data instances is heterogeneous.
+Therefore, the proposed selective fusion is more optimal for feature fusion.
+Compared with the average fusion, our selective timestep feature fusion achieves better performance by assigning the vital timestep features higher selecting weights.
+Furthermore, compared with no global-feature guidance, our proposed selective timestep fusion with global-feature guidance improves further due to the supplement of global information that enhances the ability to represent spatial distributions.
+TABLE X: OA (%) of the proposed MTMSD with different feature fusion and global-feature guidance on four datasets. the best results are shown in bold.
+Method Guidance Indian Pines PaviaU Houston 2018 Longkou
+Manual Selection 98.17 99.40 94.57 99.08
+Average Fusion 99.03 99.78 97.85 99.39
+Selective Fusion 99.31 99.90 98.07 99.53
+99.45 99.95 98.29 99.61
+IV-E Discussion and Visualization
+IV-E1 Analysis of Features from Multiple Timesteps
 To analyze the features extracted from different timestep $t$, we record the change of the classification performance when changing $t$.
 For easy understanding, we choose 4 of the 16 classes to show their changes, as illustrated in Fig.
-7.
+10.
 The performance for each class behaves differently as the $t$ increases.
@@
 For class ”Grass-pasture-mowed” and class ”corn”, features extracted at the intermediate $t$ is the most discriminative.
-Thus, an appropriate set of $t$ is vital for accurate performance.
-TABLE VII: Ablation analysis of the proposed dynamic feature fusion module with a different number of dynamic weight sets in terms of OA, AA, and $\kappa$ on the Indian Pines dataset.
-Metric The number of the sets of dynamic weights $n$
-1 2 3 4
-OA (%) 99.36 99.38 99.46 99.43
-AA (%) 99.20 99.20 99.37 99.28
-$\kappa$ 0.9918 0.9930 0.9938 0.9935
-TABLE VIII: Effect of Pretraining Steps in terms of OA, AA, and $\kappa$ on the Indian Pines dataset.
-Metric The number of pretraining steps
-0 10k 20k 30k 40k 50k
-OA (%) 68.04 97.53 98.40 99.39 99.46 99.45
-AA (%) 48.95 96.64 97.76 99.30 99.37 99.39
-$\kappa$ 58.51 0.9718 0.9817 0.9930 0.9938 0.9936
+Thus, an appropriate fusion of features at different $t$ is vital for accurate performance.
 Figure 10: Variation of classification results for four classes with the change of timestep on the Indian Pines dataset.
-TABLE IX: Classification performance with different sets of timestep $t$ on the Indian Pines dataset.
-Timestep Sets $X$ Metric
-OA (%) AA (%) $\kappa$
-$[100,200,300]$ 98.81 98.72 0.9874
-$[700,800,900]$ 98.77 98.68 0.9872
-$[100,500,900]$ 98.87 98.83 0.9886
-Dynamic Feature Fusion 99.46 99.37 0.9938
-Table IX shows the results with several sets of manually assigned $t$.
-Specifically, given a timestep set $X=[t_{1},t_{2},t_{3}]$, we extract the features at timestep $t_{1}$, $t_{2}$, and $t_{3}$ to construct the pixel-wise representation for classification.
-As Table IX presents, the performance is sensitive to the choice of $t$.
-$X$ with a large $t$, a small $t$ and a medial $t$ performs better than $X$ with only smaller $t$ or larger $t$.
-Because the former contains a wider range of $t$ leading to more diverse features which make the representation more general for samples of all classes.
-However, representation obtained by the manually designed method is not always the most appropriate for each sample, since it sacrifices some specificity to improve generalization.
-So we propose the dynamic feature fusion module to customize $T$ for each sample which makes it more accurate for classification.
-The results in Table IX demonstrate that our dynamic feature fusion module learns more effective representation compared with the manually selected method.
-Figure 8: Classification results of different patch sizes on the Indian Pines, Pavia University, and Houston2018 Dataset.
-Figure 9: Classification results of different PCA components on the Indian Pines, Pavia University, and Houston2018 Dataset.
-IV-E Parameter Analysis
-In this section, we analyze the effect of various parameters that influence classification performance by training our proposed Diff-HSI in the same experimental setting as Section IV-B with different parameters.
-IV-E1 Effect of Different Patch Size
+((a))
+((b))
+Figure 11: Feature visualization of retained and removed channels after class & timestep-oriented multi-stage purification at different timestep $t$. (a) Retained. (b) Removed.
+((a))
+((b))
+((c))
+((d))
+Figure 12: Feature visualization of different timestep $t$ with higher and lower selection weights. (a) Pseudocolor images. (b) Ground truth. (c) Higher weights. (d) Lower weights.
+((a))
+((b))
+((c))
+Figure 13: Classification results of different parameters on the Indian Pines, PaviaU, Houston 2018, and Longkou Dataset. (a) Patch size. (b) PCA component. (c) Pretraining Steps.
+IV-E2 Visualization of Multi-Timestep Multi-Stage Diffusion Feature Exploration
+The feature map of retained and removed channels after class & timestep-oriented multi-stage purification are visualized in Fig.
+11.
+The retained channels capture more semantic information compared to the removed ones after purification, as evident from Fig.
+11.
+Additionally, for the same channel, features at different timesteps exhibit diversity.
+Features at shallow timesteps tend to capture stochastic details, while those at larger timesteps focus more on higher-level semantic information.
+To further demonstrate the selective timestep feature fusion’s ability to select features at different timesteps, we visualized the features with higher and lower selection weights for the same sample.
+As shown in Fig.
+12, the results indicate that for features with higher selection weights, the response of the target pixels is more consistent with surrounding pixels of the same class, containing more classification-correlated information.
+If the features across all timesteps are averaged without discrimination, the important information relevant to the classification task may become obfuscated.
+Thus, to obtain an effective representation suitable for hyperspectral image classification, our proposed selective timestep feature fusion increases the proportion of classification-related information in the representation by assigning higher weights to more discriminative features.
+The visualization provides evidence that our proposed selective timestep feature fusion can effectively select timestep features suitable for classification tasks.
+IV-F Parameter Analysis
+In this section, we analyze the effect of various parameters that influence classification performance by training our proposed MTMSD in the same experimental setting as Section IV-B with different parameters.
+IV-F1 Effect of Different Patch Size
 First, we discuss the effect of different patch sizes on classification performance by two indexes, OA and AA.
 As shown in Fig.
-8, the patch size varies from 24$\times$24 to 60$\times$60.
+13(a), the patch size varies from 24$\times$24 to 60$\times$60.
 As the patch size increases, the performance first increases and then decreases.
@@
 Too small patches contain insufficient spatial information and too large patches reduce the attention to detailed structures.
-Thus, we choose 48$\times$48 to be the patch size for the proposed Diff-HSI.
-IV-E2 Effect of Different PCA Components
+Thus, we choose 48$\times$48 to be the patch size for the proposed MTMSD.
+IV-F2 Effect of Different PCA Components
 This section analyzes the influence of the number of PCA components, which determines how much spectral information is retained in the compressed data.
 As the number of PCA components increases, more spectral information is retained while more computational cost and more redundant information are brought.
-Since each dataset has a different number of channels, the range of PCA components is different for the three datasets.
+Since each dataset has a different number of channels, the range of PCA components is different for four datasets.
 Assuming that $N$ is the channel number of a dataset, the number of PCA components varies from $N/20$ to $N/5$.
 According to the results shown in Fig.
-9, the best performance is achieved at PCA components of $N/8$.
+13(b), the best performance is achieved at PCA components of $N/8$.
+IV-F3 Effect of Different Pretraining Steps
+We validate the effectiveness of pretraining on the four datasets in terms of OA.
+As shown in Fig.
+13(c), only 10k steps pretraining brings dramatic improvement (more than 30% OA) to the final classification performance.
+Furthermore, as the pretraining steps increase, the performance continues to rise to the best at around 40k steps.
+IV-G Efficiency Analysis
+We evaluate the inference time of different methods on four public datasets to analyze the efficiency.
+All the experiments are carried out on a Nvidia RTX 3090.
+As shown in Table XI, although our method is not the fastest in inference speed, it achieves the best classification performance since it utilizes effective multi-timestep multi-stage diffusion features.
+It is noted that our method has reduced the average inference time by 62% across four datasets compared to another diffusion-based HSI classification method, SpectralDiff.
+TABLE XI: GPU inference time(s) of different methods.
+Method Indian Pines PaviaU Houston 2018 Longkou
+2-D CNN 3.61 7.02 47.47 69.86
+3-D CNN 3.83 7.66 48.99 65.24
+SSRN 2.74 5.36 35.94 37.99
+SF 3.52 7.59 63.73 48.40
+SSFTT 3.79 7.28 31.38 59.32
+GAHT 2.51 5.25 44.90 34.19
+3DCAE 6.12 12.06 193.07 83.91
+3DAES 2.81 5.64 41.64 41.83
+UMSDFL 14.28 31.04 293.41 249.03
+SpectralDiff 10.92 22.41 268.88 178.25
+MTMSD (Ours) 3.52 12.41 168.52 61.53
 V Conclusion
 HSI contains rich spectral-spatial information and complex relations, which are critical for classification tasks.
-We propose the Diff-HSI method to learn discriminative spectral-spatial features in an unsupervised manner using the diffusion model.
-Most of the existing works perform HSI classification in an explicit way based on CNN or Transformer.
-These methods often heavily rely on the number of labeled data and are insufficient to model complex relations, limiting the improvement in performance.
-Some works explore unsupervised learning for HSI classification through an encoder-decoder feature learning method.
-However, limited by the feature-extracting capability of the model, they mainly tend to capture low-level features, lacking high-level features which are significant for discriminative spectral-spatial feature representation.
-Our proposed method leverages the powerful implicit learning ability of the diffusion model to learn effective spectral-spatial features with high degrees of freedom, constructing the step-wise feature bank.
-Furthermore, we design a dynamic feature fusion model to obtain multi-timestep representation from the feature bank automatically.
-Quantitative experiments on three HSI datasets demonstrate that our proposed Diff-HSI outperforms state-of-art supervised and unsupervised methods.
+Many supervised and unsupervised deep learning methods are proposed to learn spectral-spatial features from HSI data, achieving promising results in HSI classification.
+Recently, diffusion models as powerful models in generation and reconstruction tasks have been applied to HSI classification in one recent work.
+However, the diffusion features used in the work are extracted solely from a single timestep and a single stage of the denoising U-Net manually selected for each dataset, which limits the performance.
+Thus, we propose a diffusion-based feature learning framework that explores Multi-Timestep Multi-Stage Diffusion features for HSI classification for the first time, named MTMSD.
+Quantitative experiments on four HSI datasets demonstrate that our proposed MTMSD outperforms state-of-the-art supervised and unsupervised methods.